### v1

# AppReferences

Abstract
Principal stratification provides a causal inference framework that allows adjustment for confounded post-treatment variables when comparing treatments. Although the literature has focused mainly on binary post-treatment variables, there is a growing interest in principal stratification involving continuous post-treatment variables. However, characterizing the latent principal strata with a continuous post-treatment presents a significant challenge, which is further complicated in observational studies where the treatment is not randomized. In this paper, we introduce the Confounders-Aware SHared atoms BAyesian mixture (CASBAH), a novel approach for principal stratification with continuous post-treatment variables that can be directly applied to observational studies. CASBAH leverages a dependent Dirichlet process—utilizing shared atoms across treatment levels—to effectively control for measured confounders and facilitate information sharing between treatment groups in the identification of principal strata membership. CASBAH also offers a comprehensive quantification of uncertainty surrounding the membership of the principal strata. Through Monte Carlo simulations, we show that the proposed methodology has excellent performance in characterizing the latent principal strata and estimating the effects of treatment on post-treatment variables and outcomes. Finally, CASBAH is applied to a case study in which we estimate the causal effects of US national air quality regulations on pollution levels and health outcomes.
Keywords: Causal Inference; Dependent Dirichlet Process; Environmental Health; Principal Strata; Shared Atoms Mixture Model
1 Introduction
Principal stratification provides an effective causal inference approach to deal with post-treatment variables (Frangakis and Rubin, 2002). A post-treatment variable is defined as a variable that is potentially affected by treatment and also affects the primary potential outcomes. Thus, a post-treatment variable can be seen as an intermediate variable between the treatment and the outcome of interest. Post-treatment variables may pose inferential challenges for the estimation of treatment effects but also provide a useful approach to conceptualizing the mechanisms/pathways through which the treatment affects the outcome through the intermediate variable.
Principal stratification with respect to the post-treatment variable—also known as intermediate variable—involves categorizing subjects into different groups, known as principal strata, based on the joint potential values of that post-treatment variable at various treatment levels (Antonelli et al., 2023). Principal strata consist of units that share similar potential values of the post-treatment—which is unaffected by treatment—thus representing an inherent latent characteristic of the units.
Since the principal strata are, by definition, not influenced by the assignment of treatment, comparisons of potential outcomes at different treatment levels within a principal stratum represent well-defined causal effects (Rubin, 1974). These specific comparisons are referred to as principal causal effects. Researchers have often referred to the dissociative effect as the effect of treatment on the outcome for the stratum that includes units whose potential post-treatment variables do not notably change between different treatment levels, while labeling as the associative effect the effect of the treatment on the outcome for the stratum that includes units whose post-treatment variable do notably change between the different treatment levels (Frangakis and Rubin, 2002; Mealli and Mattei, 2012).
Most of the literature on principal stratification and principal causal effects has focused on the case of binary treatment and the binary post-treatment variable (see, e.g., Angrist et al., 1996). This implies that researchers must deal with at most four combinations of these two variables (see the broad literature for binary or discrete post-treatment variables—e.g., Cheng and Small, 2006; Imai, 2008; Ding et al., 2011; Mattei and Mealli, 2011; Mealli and Pacini, 2013; Mealli et al., 2016; Jiang et al., 2016, 2022; Mattei et al., 2024). Considering continuous post-treatment variables leads to a potentially infinite number of potential principal strata, which, in turn, makes inference and interpretation very challenging. Various solutions have been proposed in the causal inference literature to deal with continuous post-treatments. Generally, it has been proposed to: (i) dichotomize the post-treatment variable (Sjölander et al., 2009; Jiang et al., 2022), or model the continuous post-treatment variable conditional to binary treatment using (ii) fully parametric or semiparametric modeling (Jin and Rubin, 2008; Conlon et al., 2014; Lu et al., 2023), or (iii) nonparametric modeling (Schwartz et al., 2011). Recently, Antonelli et al. (2023) has proposed an extension of nonparametric modeling in the case of continuous treatment using a Bayesian methodology.
Although these methods provide potential solutions to the challenges of continuous post-treatment variables, they often require strong assumptions or subjective choices. On the one hand, dichotomization of the post-treatment variable requires the choice of the threshold on the post-treatment variable (e.g., Zigler et al., 2018; Antonelli et al., 2023; Lu et al., 2023). Although this choice might be informed by the relevant subject matter literature, it entails a level of arbitrability and requires extensive checking the robustness of the results with respect to the chosen thresholds.
The method proposed in this paper extends the exciting work of Schwartz et al. (2011). The authors model a continuous post-treatment outcomes using a Dirichlet process mixture (DPM) method (Ferguson, 1973). We build upon it to allow for a novel and more interpretable characterization of the principal strata that overcomes some challenges. Firstly, our goal is to foster better information sharing between the marginal distribution of potential post-treatment variables, so to enhance the ability to uncover the principal strata and provide a more interpretable definition of them. Secondly, while the DPM proposed by Schwartz et al. (2011) adheres to the standard stick-breaking definition by Sethuraman (1994), our approach extends this by incorporating covariates into the weights of the mixture model. This adjustment improves the imputation of missing potential outcomes for the post-treatment, consequently enhancing the estimation performance of treatment effects.
To do so, we introduce a novel Bayesian approach to principal stratification. Specifically, we introduce a Confounders-Aware SHared-atoms BAyesian mixture model (denoted by the acronym CASBAH, with the intentional alteration of the letter “H") endowed with innovative estimands for principal causal effects. CASBAH exploits the dependent Dirichlet process (Quintana et al., 2022) to establish a highly flexible distribution for the potential post-treatment variables given the confounders and the treatment variable, similar to the recent proposal of Zorzetto et al. (2024) in the context of heterogeneity in the treatment effects.
CASBAH improves the current literature on nonparametric modeling for principal stratification in many respects. Firstly, it allows for the inclusion of the confounders in the weights of the infinite Bayesian mixture model, which provides a way to control for measured confounding and a more precise estimation of the causal effects. Secondly, the proposed model—which leverages shared random kernel’s atoms—allows us to identify associative/dissociative strata borrowing information across the treatment levels. Hence, we do not have to choose subjective criteria to define the principal strata, but CASBAH can, simultaneously, (i) identify the principal strata, according to the nature of the observed data, and (ii) impute the missing post-treatment variables and outcomes. Quantification of the uncertainty of the membership of the principal strata and estimation of the principal causal effects are inherent in CASBAH.
This paper is organized as follows. In Section 2, we introduce the framework and the definition of the novel principal causal estimands. In Section 3, we propose CASBAH and how its characteristics allow us to identify the principal strata using a data-driven method. Theoretical prior properties provides insights on prior elicitation. An efficient Gibbs sampling algorithm to sample from the posterior distributions and estimate the model is also derived. Section 4 illustrates, via simulations, the ability of the proposed model to impute missing post-treatment outcomes, discover the main strata, and estimate the principal causal effects. In Section 5, we estimate the principal causal effects of air pollution policies on the mortality rate in the eastern United States (Zigler et al., 2018), where long-term fine particles pm2.5 are the post-treatment variable. In Section 6, we conclude the paper with a discussion and avenues for future work.
2 Setup
We assume that we observe $n$ independent units. For each unit $i\in\{1,\dots,n\}$, let $\mathbf{X}_{i}\in\mathcal{X}\subseteq\mathbb{R}^{P}$ the set of observed covariates, $T_{i}\in\{0,1\}$ the observed (binary) treatment, $P_{i}\in\mathcal{P}\subseteq\mathbb{R}$ the post-treatment variable, and $Y_{i}\in\mathcal{Y}\subseteq\mathbb{R}$ the response. According to Rubin’s Causal Model (Rubin, 1974), for each unit $i$, the potential outcome of the post-treatment variable is defined as $\{P_{i}(0),P_{i}(1)\}\in{\mathds{R}}^{2}$ and the potential outcome of the response is defined as $\{Y_{i}(0),Y_{i}(1)\}\in{\mathds{R}}^{2}$, for $i=1,\dots,n$. The vector $\{P_{i}(0),P_{i}(1)\}$ represents the collection of the two potential outcomes of the post-treatment variable, when the unit $i$ is assigned to the control or the treatment group, respectively. Similarly, the vector $\{Y_{i}(0),Y_{i}(1)\}$ represents the collection of the two potential outcomes for the response variable, with $Y_{i}(0)$ when the unit $i$ is assigned to the control group and $Y_{i}(1)$ when assigned to the treatment group.
2.1 Identifying Assumptions
Following the principal stratification framework (Frangakis and Rubin, 2002; VanderWeele, 2011; Mealli and Mattei, 2012; Feller et al., 2017; Ding et al., 2017; Lu et al., 2023), the principal strata are not affected by treatment assignment; therefore, a principal stratification can be used as any unit classification to define meaningful causal estimates conditional on the principal strata and discover the heterogeneous treatment effect (Mealli and Mattei, 2012).
As highlighted in the Introduction, the definition of the dissociative and associative strata is of particular interest. Formally, a dissociative stratum includes units where the treatment does not change the post-treatment variable, that is, $P_{i}(0)$ is the same as $P_{i}(1)$. In contrast, the associative stratum includes units for which treatment substantially affects the post-treatment variable, increasing it—positive associative stratum—or decreasing it—negative associative stratum.
Throughout the paper, we invoke the following assumptions:
Assumption 1 (Stable Unit Treatment Value Assumption)
$Y_{i}(T_{i})=Y_{i}(T_{1},T_{2},\cdots,T_{i},\cdots,T_{n}),\quad Y_{i}(T_{i})=Y_{i},\mbox{ for }i=1,\dots,n;$
$P_{i}(T_{i})=P_{i}(T_{1},T_{2},\cdots,T_{i},\cdots,T_{n}),\quad P_{i}(T_{i})=P_{i},\mbox{ for }i=1,\dots,n.$
The Stable Unit Treatment Value Assumption (SUTVA) enforces that each unit’s outcome is a function of its treatment only. This is a combination of (i) no interference between units, that is, the potential outcome and the potential values of the post-treatment variable of the unit $i$ do not depend on the treatment applied to other units, and (ii) consistency, that is, no different versions of the treatment levels assigned to each unit (Rubin, 1986). Under the principal stratification framework, this assumption is invoked for both the outcome variable and the post-treatment variable.
In practice, under SUTVA, for $i=1,\dots,n$, we observe only $p_{i}\in{\mathds{R}}$ and $y_{i}\in{\mathds{R}}$, that is, the realization of the random variable $P_{i}$ and $Y_{i}$, respectively, and defined as $P_{i}=(1-T_{i})\cdot P_{i}(0)+T_{i}\cdot P_{i}(1)$ and $Y_{i}=(1-T_{i})\cdot Y_{i}(0)+T_{i}\cdot Y_{i}(1)$.
Assumption 2 (Strongly Ignorable Treatment Assignment)
Given the observed covariates ${x}_{i}$,
$\{\bigl{(}Y_{i}(1),Y_{i}(0)\bigl{)},\bigl{(}P_{i}(0),P_{i}(1)\bigl{)}\}\perp\!\!\!\perp T_{i}\mid X_{i},$
$0<Pr\left(T_{i}=1\mid X_{i}=x\right)<1\>\>\forall\>x\in\mathcal{X},$
where $\mathcal{X}$ is the features’ space.
This assumption states that: (i) for each unit, the potential outcome and the potential values of the post-treatment variable are independent of the treatment conditional on the set of covariates $X_{i}$; (ii) all units have a positive chance of receiving the treatment.
3 Confounders-Aware Shared-atoms Bayesian Mixture Model
Following the Bayesian paradigm and invoking the De Finetti’s theorem (Rubin, 1978), the joint probability distribution of the involved variables is unit exchangeable and it is defined as
$\Pr(Y(0),Y(1),P(0),P(1),T,X)=\int_{\Theta}\prod_{i=1}^{n}\Pr(Y_{i}(0),Y_{i}(1),P_{i}(0),P_{i}(1),T_{i},X_{i},\theta)p(\theta)d\theta,$
where $p(\theta)$ is the prior distribution for all the involved parameters $\theta$ that take values in the parameter space $\Theta$. The inner probability can be factorized as
$\Pr(T_{i}\mid Y_{i}(0),Y_{i}(1),P_{i}(0),P_{i}(1),X_{i},\theta)$$\times\Pr(Y_{i}(0),Y_{i}(1)\mid P_{i}(0),P_{i}(1),X_{i},\theta)$
×$\times\Pr(P_{i}(0),P_{i}(1)\mid X_{i},\theta)$$\times\Pr(X_{i}\mid\theta).$ (1)
Following the assumptions in the previous section, the conditional probability of the treatment variable can be written as $\Pr(T_{i}\mid Y_{i}(0),Y_{i}(1),P_{i}(0),P_{i}(1),X_{i},\theta)=\Pr(T_{i}\mid X_{i},\theta)$. Moreover, we condition the empirical distribution of covariates so that $\Pr(X_{i}\mid\theta)=\Pr(X_{i})$.
However, the remaining two probabilities in equation (1), have to be modeled: (i) the distribution of potential outcomes of the response conditional on the potential outcome of the post-treatment variable and covariates, and (ii) the distribution of the potential outcome of the post-treatment variable conditional on the covariates.
We focus our attention on the latter, that is, the distribution of the potential outcome of the post-treatment variable conditional on the covariates for which we propose a novel Bayesian nonparametric approach. As our primary focus lies in the post-treatment variable distribution, for the sake of clarity in our discussion, we make the simplifying assumption of employing a more straightforward parametric model for the conditional response distribution following the settings of Schwartz et al. (2011). Although the outcome model is assumed to be a linear regression model, it can also be generalized to a more complex and flexible model.
For the post-treatment variable, we previously exploited a dependent nonparametric mixture, following the approach adopted in the previous section of this paper and more broadly the dependent Dirichlet process (MacEachern, 2000; Barrientos et al., 2012; Quintana et al., 2022). In particular, our model also shares some similarities with the hierarchical Dirichlet process (Teh et al., 2004, 2006; Teh and Jordan, 2009) and the more recent common atom model of Denti et al. (2023).
Specifically, we assume for each $i=1,\dots,n$:
$\{P_{i}\mid{x}_{i},t\}$ $\sim f^{(t)}(\cdot\mid{x}_{i}),\mbox{ for }t=\{0,1\},$
[…]
$\{G_{{x}_{i}}^{(0)},G_{{x}_{i}}^{(1)}\}$ $\sim\Pi,$ (2)
where ${\mathcal{K}}(\cdot;x,\psi)$ is a continuous density function, for every $\psi\in\Psi$, and $G_{{x}_{i}}^{(t)}$ is a random probability measure depending on the confounders $x_{i}$ associated to an observation assigned to treatment level $t$. In particular, we have two treatment levels $t=\{0,1\}$, therefore, the distributions belonging to $G_{{x}_{i}}^{(0)}$ and $G_{{x}_{i}}^{(1)}$ are characterized by specific weights assigned to a common set of atoms induced by the common process $\Pi$.
Following the characterization of the random probability measures we can write:
$G_{{x}_{i}}^{(t)}=\sum_{l\geq 1}\pi_{l}^{(t)}({x}_{i})\delta_{\mathbb{\psi}_{l}},$ (3)
where the sequences $\{\pi_{l}^{(t)}(x_{i})\}_{l\geq 1}$ for $t=\{0,1\}$ represent infinite sequences of random weights, and $\{\psi_{l}\}_{l\geq 1}$ is an infinite sequence of random kernel’s parameters—independent and identically distributed from a base measure $H$—shared among potential outcome of post-treatment variable under both the treatment level $t$.
Furthermore, the sequences of weights are defined through a stick-breaking representation (Sethuraman, 1994),
$\pi_{l}^{(t)}({x}_{i})$ $=\omega_{l}^{(t)}({x}_{i})\prod_{r<l}\{1-\omega_{r}^{(t)}({x}_{i})\};$ (4)
where $\{\omega_{l}^{(t)}({x})\}_{l\geq 1}$ , for $t=\{0,1\}$ are $[0,1]$-valued independent stochastic processes.
The discrete nature of the random probability measure $G_{{x}_{i}}^{(t)}$, for $t=\{0,1\}$, allows us to introduce the latent categorical variables $S_{i}^{(t)}$, for $t=\{0,1\}$, describing the groups of units defined by heterogeneity in $P_{i}(t)$. Assuming $\mathds{P}\{S_{i}^{(t)}=l\}=\pi_{l}^{(t)}({x}_{i})$, we can write model in (2), exploiting conditioning on $S_{i}^{(t)}$, as
$\{P_{i}|{x}_{i},t,\psi,S_{i}^{(t)}=l\}\sim{\mathcal{K}}(\cdot\mid{x}_{i},\psi_{l}),\quad\psi_{l}\sim H.$
where $\psi$ represents the infinite sequence $\{\psi_{l}\}_{l\geq 1}$.
Among the plethora of dependent non-parametric processes (see the recent review by Quintana et al., 2022, for a detailed description), we focused on the dependent probit stick breaking process for its success in applications, its good theoretical properties and ease of computation (Rodriguez and Dunson, 2011). Consistently with this, we specify:
$\omega_{l}^{(t)}(x_{i})=\Phi(\alpha_{l}^{(t)}(x_{i})),\quad\alpha_{l}^{(t)}(x_{i})\sim\mathcal{N}(\beta_{l0}^{(t)}+x_{i}^{T}\beta_{l}^{(t)},1),$ (5)
where $\Phi(\cdot)$ is the CDF of a standard Normal distribution, $\{\alpha_{l}^{(t)}(x_{i})\}_{l\geq 1}$ has Gaussian distributions with mean a linear combination of the $p$ covariates $x_{i}$, and $\{\beta_{l0}^{(t)},\beta_{l}^{(t)}\}_{l\geq 1}$ are the probit regression parameters treatment-specific.
As commonly done, we assume the kernel $\mathcal{K}$ to be a Gaussian, so that model (2)–(3) specifies to
$\{P_{i}(t)|{x}_{i},S_{i}^{(t)}=l,\eta,\sigma\}\sim\mathcal{N}(\eta_{l},\sigma_{l}^{2}).$ (6)
where $\eta$ and $\sigma$ represent infinite sequences of location parameters $\{\eta_{l}\}_{l\geq 1}$ and scale parameters $\{\sigma_{l}\}_{l\geq 1}$, respectively, such that $\psi_{l}=(\eta_{l},\sigma_{l})$.
Prior elicitation is completed by assuming for the parameters $\eta_{l}$ and $\sigma_{l}$ in (6)
$\eta_{l}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(\mu_{\eta},\sigma_{\eta}^{2}),\mbox{ and }\sigma_{l}^{2}\stackrel{{\scriptstyle iid}}{{\sim}}\mbox{InvGamma}(\gamma_{1},\gamma_{2}).$
where InvGamma($\gamma_{1},\gamma_{2}$) represents the inverse gamma distribution with the shape parameter $\gamma_{1}\in\mathbb{R}^{+}$ and the scale parameter $\gamma_{2}\in\mathbb{R}^{+}$, and mean equal to $\frac{\gamma_{2}}{\gamma_{1}-1}$ and variance $\frac{\gamma_{2}^{2}}{(\gamma_{1}-1)^{2}(\gamma_{1}-2)}$.
For the regression parameters in (5), we assume the following multivariate Gaussian prior:
$\beta^{(t)}\sim\mathcal{N}_{(p+1)(L-1)}(\xi,\Omega)$ (7)
for $t=0,1$ and $l\geq 1$. According to Fasano et al. (2022), the Gaussian prior leads to a straightforward posterior computation, as discussed in the next section.
3.1 Causal Estimands
In this paper, we introduce a formal and novel definition of principal strata that takes into account the continuous nature of the post-treatment variable.
We assume that the units can be marginally—respect to the treatment level—grouped into clusters with finite first moments. Consequently, each unit $i$ belongs to either: the dissociative stratum if $S_{i}(1)=S_{i}(0)$, which is equivalent to $\mathbb{E}_{S_{i}^{(1)}}[P_{i}(1)]=\mathbb{E}_{S_{i}^{(0)}}[P_{i}(0)]$; or the associative positive stratum if $S_{i}(1)\not=S_{i}(0)$ and $\mathbb{E}_{S_{i}^{(1)}}[P_{i}(1)]>\mathbb{E}_{S_{i}^{(0)}}[P_{i}(0)]$ (now on, we indicate these conditions as $S_{i}(1)\succ S_{i}(0)$); or the associative negative stratum if $S_{i}(1)\not=S_{i}(0)$ and $\mathbb{E}_{S_{i}^{(1)}}[P_{i}(1)]<\mathbb{E}_{S_{i}^{(0)}}[P_{i}(0)]$ (now on, we indicate these conditions as $S_{i}(1)\prec S_{i}(0)$).
Definition 1 (Causal Estimands)
Conditionally on strata allocation, we can define the causal estimands for the expected dissociative effect (EDE) and the expected associative effects (EAE) as
$EDE=\tau_{0}=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid S_{i}(1)=S_{i}(0)],$
$EAE_{+}=\tau_{+1}=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid S_{i}(1)\succ S_{i}(0)],$
$EAE_{-}=\tau_{-1}=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid S_{i}(1)\prec S_{i}(0)].$
These causal estimands are not directly identifiable from the data due to the conditioning on unobservable quantities (i.e., $S_{i}(1),S_{i}(0)$). However, they can be weak identified invoking the causal assumptions introduced above—i.e., SUTVA and strongly ignorable treatment assignment.
Theorem 1
Invoking assumptions (1) and (2), the expected value of each potential outcome can be rewritten as the following
$\mathbb{E}\big{[}Y_{i}(t)\mid S_{i}(1)=s_{1},S_{i}(0)=s_{0},X_{i}=x\big{]}$
$\quad=\int_{p_{0}p_{1}}\mathbb{E}\big{[}Y_{i}\mid T_{i}=t,X_{i}=x,P_{i}(1)=p_{1},P_{i}(0)=p_{0}\big{]}$
×$\quad\quad\times Pr\big{(}P_{i}(1)=p_{1},P_{i}(0)=p_{0}\mid T_{i}=t,S_{i}(1)=s_{1},S_{i}(0)=s_{0},X_{i}=x\big{)}dp_{0}p_{1}dx;$ (8)
where the inner expectations $\mathbb{E}\big{[}Y_{i}\mid T_{i}=t,X_{i}=x,P_{i}(1)=p_{1},P_{i}(0)=p_{0}\big{]}$ is estimate with the outcome model $Y_{i}\mid T_{i},X_{i},P_{i}(1),P_{i}(0)$, as well the probability $\Pr\big{(}P_{i}(1)=p_{1},P_{i}(0)=p_{0}\mid T_{i}=t,S_{i}(1)=s_{1},S_{i}(0)=s_{0},X_{i}=x\big{)}$, that is defined by the model for the potential post-treatment variables.
The proof is reported in Appendix B.
3.2 Discovery of Principal Strata
The definition of the previous (3) for the post-treatment variable model allows us to characterize two distinct specific mixtures for each of the two treatment levels $t=\{0,1\}$, taking into account the different heterogeneity in the potential post-treatment variables induced by the covariates $X$. Simultaneously, the parameters of the shared random kernel $\{\psi_{l}\}_{l\geq 1}$ induce desirable characteristics in the potential post-treatment variables, such as (i) the potential post-treatment variables $\{P(1),P(0)\}$ are not independent, even conditionally, of the covariates $X$—this is clear from the posterior distribution reported in Section 3.3 and Appendix D—and (ii) there is a nonnull probability that any $i$ belongs to the defined dissociative stratum—see Theorem 2.
The fundamental element in explaining this is the definition of the latent categorical variables $\{S_{i}^{(0)},S_{i}^{(1)}\}$ for each unit $i\in\{1,\dots,n\}$, which describe the probability that each unit will be allocated in the components of the mixture. The Bayesian nonparametric mixtures are well-known for their ability to capture and describe the cluster allocation of the observations; however, we focus on the causal inference goal to describe the principal strata.
In our fully Bayesian approach, pairs $\{S_{i}^{(0)},S_{i}^{(1)}\}$ are random variables that define the two marginal distributions of the random unit partition (Quintana, 2006), based on the two potential post-treatment variables. However, we want to define the join distribution of them, specifically the unit allocation in the principal strata. This join distribution is not directly observed from the data, due to missing data problems, but we can estimate this thanks to the characteristic structure of the proposed CASBAH model.
The proposed CASBAH model is defined such that the atoms $\{\psi_{l}\}_{l\geq 1}$ are shared between the two potential outcomes of the post-treatment variable of the same unit. Therefore, the model informs us about the posterior probability that $\{S_{i}^{(0)},S_{i}^{(1)}\}$ is associated with the same or different atoms, for each unit $i$. For example, the probability for the unit $i\in\{1,\dots,n\}$ to belong to the dissociative stratum—i.e., $S_{i}(0)=S_{i}(1)$—is equivalent to the probability that the atom for the control is the same for the treated. Such a probability is not null and can be derived, a priori, by means of the following theorem.
Theorem 2
Given a unit $i\in\{1,\dots,n\}$, with covariates $x_{i}$, following the (2) distribution, the probability to belong to the dissociative stratum for each unit $i$ is a function of its covariates $x_{i}$, defined as following
$\Pr[S_{i}(0)=S_{i}(1)]=\frac{\rho_{2}(x_{i})[(1+\rho_{2}(x_{i})-2\rho_{1}(x_{i}))^{L}-1]}{\rho_{2}(x_{i})-2\rho_{1}(x_{i})}$ (9)
where $\rho_{1}(x_{i})=\mathbb{E}\big{[}\Phi\big{(}\alpha(x_{i})\big{)}\big{]}$ and $\rho_{2}(x_{i})=\mathbb{E}\big{[}\Phi\big{(}\alpha(x_{i})\big{)}^{2}\big{]}$ and $0\leq\Pr[S_{i}(0)=S_{i}(1)]\leq 1$.
The proof is reported in Appendix C.
[…]
Figure 1: Probability to belong to the dissociative stratum. Function varies according to different values of $\alpha(x)=\beta_{0}+\beta_{1}x$.
3.3 Posterior Inference
As Theorem 1 shows, the principal causal estimand are functions of the models for potential outcomes and the potential post-treatment variable that involve parameters in both models. Following the Bayesian paradigm, the estimation is based on the posterior distribution of the parameters conditional on the data.
Sampling from the posterior joint distribution is straightforward via Gibbs sampling. In particular, the algorithm—described in detail in Section D—takes inspiration from those proposed by Teh et al. (2004, 2006); Teh and Jordan (2009), for the hierarchical structure, and by Fasano et al. (2022) for the probit regression in the weights.
Assuming a general linear regression for the outcome model that depends on the parameter $\theta_{0}^{Y}$ for the outcome under control and $\theta_{1}^{Y}$ for the treated outcome, the Gibbs sampler interests the following distributions:
  • The latent variable $S_{i}^{(t)}$ is a multivariate distribution with
    $\Pr\{S_{i}^{(t)}=l\}\propto\pi_{l}^{(t)}({x}_{i})\Pr(P_{i}(t)\mid S_{i}^{(t)}=l,X_{i};\eta_{l},\sigma_{l}),$
    where $\pi_{l}^{(t)}({x}_{i})=\Phi(\alpha_{l}^{(t)}({x}_{i}))\prod_{r<l}(1-\Phi(\alpha_{r}^{(t)}({x}_{i})))p$ for $l=1,\dots,L-1$ and $\Phi(\alpha_{L}^{(t)}({x}_{i}))=1$.
  • The cluster-specific parameters are drawn from
    $\psi_{l}\propto\Pi_{x}\prod_{i:S_{i}^{(0)}=l}\Pr(P_{i}(0)\mid S_{i}^{(0)}=l,X_{i};\psi_{l})\prod_{i:S_{i}^{(1)}=l}\Pr(P_{i}(1)\mid S_{i}^{(1)}=l,X_{i};\psi_{l}).$
  • The logit regression parameters for conditional-dependent weights are drawn from
    $\beta^{(t)}\propto\Pr(\beta^{(t)})\prod_{i=1}^{n}\Pr(P_{i}(t)\mid S_{i}^{(t)},X_{i};\beta^{(t)})$
    following (Fasano et al., 2022) demonstration—detail in Appendix A.
  • The missing post-treatment variable is imputed from
    $P_{i}(1-T_{i})\propto\Pr(Y_{i}\mid P_{i}(0),P_{i}(1),X_{i};\theta_{0}^{Y})^{1-T_{i}}\Pr(Y_{i}\mid P_{i}(0),P_{i}(1),X_{i};\theta_{1}^{Y})^{T_{i}}\Pr(P_{i}(1-T_{i})\mid X_{i};\psi).$
  • The posterior distribution for the outcome model is built from
    $\theta_{t}^{Y}\sim\Pr(\theta_{t}^{Y})\prod_{i=1}^{n}\text{P}(Y_{i}(t)\mid P_{i}(0),P_{i}(1),X_{i};\theta_{t}^{Y}).$
4 Simulation Study
The performance of the proposed CASBAH mixture model is evaluated by means of a simulation study. Our objective is to investigate the model’s ability to (i) accurately impute the missing post-treatment and outcome variables, that is, to assess the potential bias of the expected value $E[P_{i}(1)-P_{i}(0)]$ and $E[Y_{i}(1)-Y_{i}(0)]$ over the units $i=1\dots,n$—, (ii) correctly identify the principal strata. To achieve this, we conducted simulations under five different data-generating mechanisms and analyzed the results to understand the model’s behavior in different scenarios.
Specifically, we assume a linear regression model for the outcome model, defined as follows:
$\begin{bmatrix}Y(0)\\ Y(1)\end{bmatrix}\sim\mathcal{N}_{2}\begin{pmatrix}\begin{bmatrix}\theta_{00}+\theta_{01}P(0)\\ \theta_{10}+\theta_{11}P(1)+\theta_{12}P(0)+\theta_{13}P(0)P(1)\end{bmatrix},\begin{bmatrix}e^{\lambda_{0}}&0\\ 0&e^{\lambda_{0}+\lambda_{1}P(1)}\end{bmatrix}\end{pmatrix}.$ (10)
Assuming as prior distribution for the parameters
[…]
with $\mu_{\theta}=0$, $\sigma_{\theta}=10$, $\mu_{\lambda}=0$, and $\sigma_{\lambda}=2$. Clearly, our model can accommodate more complex choices for the outcome, but we decided to opt for a simpler model following what had already been done in the literature by Schwartz et al. (2011). The performance of the proposed approach is compared to those obtained with the Schwartz et al. (2011)’s model (hereafter referred to as SLM).
For the data generating process, we simulate two Bernoulli confounders $(X_{1},X_{2})$ for scenarios 1 to 4 and five Bernoulli confounders $(X_{1},X_{2},X_{3},X_{4},X_{5})$ for scenario 5, and a binary treatment variable such that $T_{i}\sim\mbox{Be}(\mbox{logit}(0.4X_{1i}+0.6X_{2i}))$ for the scenarios $1-4$ and $T_{i}\sim\mbox{Be}(\mbox{logit}(0.4X_{1i}+0.6X_{2i}-0.3X_{3i}+0.2X_{4i}X_{5i}))$ for scenario 5, for $i=1,\dots,n$.
Each scenario assumes a different conformation of the strata for the continuous post-treatment variable $P_{i}=(P_{i}(0),P_{i}(1))\in\mathbb{R}^{2}$. Each stratum is obtained by introducing, for each unit, categorical variables $S_{i}(0)$ and $S_{i}(1)$, for control and treatment levels, respectively, with the vector of probabilities that depend on the values of the confounders and allocates the unit in different clusters. Conditionally on the cluster allocation $S_{i}^{(t)}=s$, with $s$ that has the same support for control and treatment levels, we simulate both potential post-treatment variables—under control and under treatment, respectively—as
$P_{i}(0)|S_{i}(0)=s\sim\mathcal{N}(\eta_{s},\sigma^{2}_{s}),\quad P_{i}(1)|S_{i}(1)=s\sim\mathcal{N}(\eta_{s},\sigma^{2}_{s}).$
Continuous potential outcomes $(Y_{i}(0),Y_{i}(1))$, for $i=1,\dots,n$, are simulated following the model (10) with common values for $\lambda_{0}=-0.5$ and $\lambda_{1}=0.1$, while $\theta^{(0)}$ and $\theta^{(1)}$ are different for each scenario. In each setting, the sample size is set to $n=500$. For each scenario, we simulate $100$ samples. Below we provide further details for each one of the five scenarios.
Scenario 1: We investigate a situation in which there are two strata: one with a dissociative effect and one with a positive associative effect. In particular, for $S_{i}(0)=S_{i}(1)=1$ $P_{i}(0),P_{i}(1)\sim\mathcal{N}(1,0.05)$, and for $S_{i}(0)=2$ and $S_{i}(1)=3$ $P_{i}(0)\sim\mathcal{N}(2,0.05)$ and $P_{i}(1)\sim\mathcal{N}(3,0.05)$. The regression parameters for the Y-model are $\theta^{(0)}=(1,2)$ and $\theta^{(1)}=(1,2,-1,0.5)$.
Scenario 2: We focus on a case where we have a dissociative stratum—for $S_{i}(0)=S_{i}(1)=1$ both $P(0)$ and $P(1)$ are simulated from $\mathcal{N}(2,0.05)$—, an associative stratum with a positive effect—for $S_{i}(0)=2$ and $S_{i}(1)=3$ $P_{i}(0)\sim\mathcal{N}(2,0.05)$ and $P_{i}(1)\sim\mathcal{N}(3,0.05)$—, and dissociative stratum with a negative effect—for $S_{i}(0)=2$ and $S_{i}(1)=1$ $P_{i}(0)\sim\mathcal{N}(2,0.05)$ and $P_{i}(1)\sim\mathcal{N}(1,0.05)$. The regression parameters for Y-model are $\theta^{(0)}=(1,2)$ and $\theta^{(1)}=(1,1.2,-1,1)$.
Scenario 3: This scenario corresponds to Scenario 1 with closer atoms for the strata and different variances. In particular, the dissociative stratum has $(P_{i}(0)|S_{i}(0)=1)=(P_{i}(1)|S_{i}(1)=1)\sim\mathcal{N}(1.5,0.12)$, and the associative stratum has $(P_{i}(0)|S_{i}(0)=2)\sim\mathcal{N}(2,0.1)$ and $(P_{i}(1)|S_{i}(1)=3)\sim\mathcal{N}(2.5,0.08)$. The regression parameters for Y-model are $\theta^{(0)}=(1,2)$ and $\theta^{(1)}=(1,1.2,-0.8,0.5)$.
Scenario 4: This scenario corresponds to Scenario 2 with closer atoms for the strata and different variances. In particular, the dissociative stratum has $(P_{i}(0)|S_{i}(0)=1)=(P_{i}(1)|S_{i}(1)=1)\sim\mathcal{N}(1.5,0.12)$, the associative positive stratum has $(P_{i}(0)|S_{i}(0)=2)\sim\mathcal{N}(2,0.1)$ and $(P_{i}(1)|S_{i}(1)=3)\sim\mathcal{N}(2.5,0.08)$, and the associative negative stratum has $(P_{i}(0)|S_{i}(0)=2)\sim\mathcal{N}(2,0.1)$ and $(P_{i}(1)|S_{i}(1)=1)\sim\mathcal{N}(1.5,0.12)$. The regression parameters for Y-model are $\theta^{(0)}=(1,2)$ and $\theta^{(1)}=(1,1.2,-0.8,0.5)$.
Scenario 5: We investigate the scenario with the three strata when the number of confounders increases, in particular the treatment variable and cluster allocation variables that depend on five confounders. The dissociative stratum has $(P_{i}(0)|S_{i}(0)=1)=(P_{i}(1)|S_{i}(1)=1)\sim\mathcal{N}(2,0.05)$, the associative positive stratum has $(P_{i}(0)|S_{i}(0)=3)\sim\mathcal{N}(3,0.05)$ and $(P_{i}(1)|S_{i}(1)=4)\sim\mathcal{N}(4,0.05)$, and the associative negative stratum has $(P_{i}(0)|S_{i}(0)=2)\sim\mathcal{N}(2,0.05)$ and $(P_{i}(1)|S_{i}(1)=1)\sim\mathcal{N}(1,0.05)$. The regression parameters for Y-model are $\theta^{(0)}=(1,2)$ and $\theta^{(1)}=(1,1.2,-1,0.5)$.
We choose the same hyperparameters for each setting such that the prior is non-informative and in common for all the settings. For the regression parameters in (5) and for the parameters $\eta_{l}$ and $\sigma_{l}$ in (6) we use the following conjugate priors
$\beta^{(t)}\sim\mathcal{N}_{(p+1)(L-1)}(0,20\times\mathbf{1}_{(p+1)(L-1)}),$
$\eta_{l}\sim\mathcal{N}(0,20),\mbox{ and }\sigma_{l}\sim\mbox{InvGamma}(2,0.5),$
for $t\in\{0,1\}$, $l\in\{1,\dots,20\}$, and $p$ according with the covariates considered in different settings, and where $\mathbf{1}_{q}$ is a diagonal matrix $q\times q$.
Table 1 reports the median and interquartile range (IQR) of the bias for the expected value of the posterior distribution of sample average of $P_{i}(1)-P_{i}(0)$ and $Y_{i}(1)-Y_{i}(0)$, for $i\in\{1,\dots,n\}$. The results for the five scenarios for the proposed model CASBAH show a good ability of the proposed model to impute the missing variables and capture the true distribution, indeed the medians of the bias are close to zero and the interquartile range is reasonable according to the simulated variability. The comparison with the Schwartz et al. (2011)’s model, SLM, underlines the superiority of CASBAH, which obtains medians of the bias closer to zero and a more contained IQR for $\mathbb{E}[P_{i}(1)-P_{i}(0)]$ in all scenarios and almost of them for $\mathbb{E}[Y_{i}(1)-Y_{i}(0)]$.
Scenario 1 Scenario 2 Scenario 3 Scenario 4 Scenario 5
Post-treatment variable
CASBAH Median -0.0164 0.0027 -0.0318 0.0027 0.0011
IQR 0.0315 0.0254 0.0462 0.0325 0.0183
SLM Median -0.0637 0.3388 0.1254 0.3280 0.3572
IQR 0.0989 0.0986 0.0956 0.0963 0.1216
Outcome
CASBAH Median 0.0024 0.0256 0.0067 0.0419 0.0344
IQR 0.2235 0.1550 0.4661 0.1357 0.1316
SLM Median -1.9891 -1.8265 -1.4363 -1.3115 -1.8856
IQR 0.2546 0.2110 0.1904 0.1836 0.3727
Table 1: Median and interquartile range (IQR) of the bias for the expected value of the posterior distribution of sample average of $P_{i}(1)-P_{i}(0)$ and $Y_{i}(1)-Y_{i}(0)$, for $i\in\{1,\dots,n\}$. Values reported for CASBAH and SLM.
To evaluate the correct identification of the principal strata, we use the adjusted Rand index (ARI, introduced by Hubert and Arabie (1985)). The ARI values for the five scenarios are reported in Table 2. For all scenarios, the index is close to $1$ and confirms that the proposed CASBAH model can correctly identify the principal strata and, combined with the good missing data imputation, allows us to estimate the expected associative and dissociative effects. The SLM model does not identify the principal strata according to our definition or simultaneously to the model estimation, in opposition to CASBAH.
Scenario 1 Scenario 2 Scenario 3 Scenario 4 Scenario 5
Mean 0.9850 0.9906 0.9706 0.9717 0.9154
SD 0.0997 0.0597 0.1021 0.0892 0.1577
Table 2: Adjusted rand index for the five simulated scenarios computed on the point estimated partitions obtained with the proposed model CASBAH. mean and empirical standard deviation (sd) are reported.
5 Assessing the Impact of Environmental Policies on Particulate Pollution and Health
As recently as 7 February 2024 the US Environmental Protection Agency (EPA) has proposed a revision of the National Ambient Air Quality Standards (NAAQS) (U.S. Environmental Protection Agency, 2024), which will result in a more stringent environmental policy for the concentration of air pollutants in the US that cause health hazards. To gather information and better understand the implementation of this revised policy, it is critically important to evaluate how past air quality regulations benefitted health outcomes and how these effects varied between different communities.
In our application, we explore this relevant research question by investigating the repercussions of a previous revision of NAAQS, leveraging insights gleaned from historical data and health data. First, we evaluate the direct and indirect effects of the 2005 NAAQS revision on the mortality rate in the period 2010-2016, considering the variation of levels pm2.5, under the principal stratification framework. Second, we take advantage of the flexible Bayesian non-parametric mixture model for the post-treatment variable (that is, pm2.5) to understand how these effects can vary within strata. Third, we characterize the different identified groups of individuals.
5.1 Data and Study Design
To answer our research question, we have merged two datasets: (i) one dataset containing the information about the NAAQS designations, as well as air pollution levels and the demographic and socioeconomic characteristics in the counties in the eastern U.S., used in the Zigler et al. (2018)’s analysis (data can be found at Zigler, 2017), and (ii) one dataset containing the information on the age-adjusted mortality rate in these counties available by the Center for Disease Control and Prevention of U.S. (Friede et al., 1993).
The initial dataset from Zigler (2017) is made up of 482 counties in the eastern U.S., where national monitoring networks have detected the concentration of pm2.5. In 2005, the EPA designated as non-attainment of the NAAQS these counties where the average concentration of pm2.5 was above $15\mu g/m^{3}$, or otherwise attainment (note that the revision occurred in 1997, but became effective just in 2005 due to legal disputes). States containing counties designated as non-attainment were required to develop or revise State Implementation Plans (SIP) outlining how a non-attainment area will achieve standards with strategies to reduce ambient concentrations of pm2.5.
As noted in the previous literature, there is a large diversity in local actions in response to the non-attainment designation and in the enactment of subsequent SIPs (Greenstone, 2004; Zigler et al., 2018). Due to this diversity, the direct analysis of the effect of the designation is very similar to an intention-to-treat analysis. This highlights the importance of conducting a principal stratification analysis to compare the effects on health outcomes across different locations: specifically, it contrasts the effects in areas where pollution was effectively reduced by the non-attainment designation (associative negative effects) with those areas where pollution was not measurably affected (dissociative effect), or where it actually increased (associative positive effects).
The study design specifies a baseline period from 2000 to 2005. During this time, each county, $i$, in the dataset is classified according to its attainment status under the EPA 2005 NAAQS revisions. We define a binary treatment variable, $T_{i}$, where $T_{i}=1$ indicates that the i-th county was designated as non-attainment and therefore was required to develop or revise its State Implementation Plans. In contrast, $T_{i}=0$ indicates that the i-th county was designated as attainment and did not have these requirements.
For this period, we also have the average ambient concentration of pm2.5(from the EPA monitoring locations within each county), as well as a number of counfounders such as census variables like: the percentage of Hispanic and black residents; the average household income; the percentage of females; the average house value; the proportion of residents in poverty; the proportion of residents with a high school diploma; the smoking rate; the population; the percentage of residents in urban area; the employment rate (the percentage of the workforce employed); the percentage of the move in the last 5 years. We also have meteorological variables such as the averages of daily temperatures and the relative average of humidity; the dew point (the temperature at which air becomes saturated with moisture, leading to the formation of dew or condensation).
Consistently with Zigler et al. (2018), we identify as the follow-up period a period after 5 years from the implementation of the 2005 NAAQS revision. For this period, we have data on the levels of pm2.5(2010-2012) from Zigler (2017). The levels of pm2.5in the follow-up period are compared with the levels in the baseline period to establish the decrease / increase in air pollution following the implementation of NAAQS. This variable serves as the post-treatment variable $P_{i}$. Furthermore, we gather the publicly available age-adjusted mortality rate of all-cause mortality (2010-2016) from the website of the US Center for Disease Control and Prevention (CDC) (Friede et al., 1993). Mortality rates in the follow-up period are also contrasted with the rates in the baseline period to assess the impact of NAAQS on mortality rates accounting for baseline rates. This variable serves as our outcome variable $Y_{i}$.
The final dataset, obtained by merging the previously mentioned data sources and cleaning the data set to avoid missing data, comprises 384 counties, 270 of which were designed as attainment and 114 as non-attainment. In Appendix, Figure F.1 maps this final analysis data set.
5.2 Results
We apply the proposed CASBAH model to the 384 counties in the eastern US described above, including all covariates —- census and meteorological variables —- in the weights of the post-treatment variable mixture, while for the outcome model we use the linear model in (10). The model identifies the three strata: the dissociative stratum with 115 counties ($30\%$ of the total counties analyzed), the associative positive stratum with 48 counties ($12.5\%$), and the associative negative stratum with 221 counties ($57.5\%$). According to the definition of the three strata and as visualized in the image on the left in Figure 2, the dissociative stratum—identified with the color yellow—is composed of counties where the NAAQS revision does not substantially affect the level of pm2.5 , in fact, the expected value of $\mathbb{E}[P_{i}(1)-P_{i}(0)|S_{i}(1)=S_{i}(0)]$ for the counties allocated to this strata has a median close to zero and the first and third quantiles of $-0.29\mu g/m^{3}$ and $0.07\mu g/m^{3}$. The associative negative stratum—identified with the color green—is made up of counties where the implementation of environmental plans significantly decreases the levels of pm2.5 . Specifically, in these counties, the NAAQS revision reduced by $0.9\mu g/m^{3}$ the median levels pm2.5 , with first and third quantiles of $-1.46\mu g/m^{3}$ and $-0.38\mu g/m^{3}$. The associative positive stratum—identified with the color red—is made up of counties where the revision of NAAQS was associated with increases in pm2.5 levels by $0.8\mu g/m^{3}$ in the median. However, this latter stratum has the highest uncertainty for $\mathbb{E}[P_{i}(1)-P_{i}(0)|S_{i}(1)\prec S_{i}(0)]$, with values of the first and third quantiles equal to $0.13\mu g/m^{3}$ and $1.34\mu g/m^{3}$, respectively.
The corresponding distributions of the expected dissociative/associative effects are reported in the right image in Figure 2 and show the effect of the attainment or non-attainment designations on the mortality rate conditional to the three strata—i.e., conditional to the heterogeneity in the causal effects of the NAAQS revision on the level of pollution. The EDE and the EAE+ have a similar increase in the median $2.7{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$ of the mortality rate when environmental plans are applied. The means of these two effects are $0.8{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$ and $1.3{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$, respectively. For both principal causal effects, the interquartile interval includes the zero value. This means that in counties where the NAAQS revision does not affect the level pm2.5 or where this level increases, the mortality rate is also not affected. The EAE-—in green color on the right image of Figure 2—assumes negative values, indicating that the implementation of environmental plans, in the counties where the regulations affect the level of pm2.5 , significantly reducing it, also reduces the median age adjusted mortality rate. Specifically, the mortality rate decreases by $11.9{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$ in the median and with an interquartile interval of $[-15.2{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}},-9.1{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}]$.
Figure 2: Boxplots of the three identified strata for (left) the conditional average of the difference of the post-treatment variables, and (right) the expected associative/dissociative effects. The light-blue vertical lines show the value zero, identifying the null effects.
In addition, it is our interest to characterize the three different strata (Figure F.3 in the Supplementary Materials visualizes the average of observed covariates within each stratum—reported as colored lines—and compares them with the average of covariates between the full 384 counties). The associative positive stratum and the dissociative stratum are composed of rural counties with a smaller population density, where the population has lower income with respect to the mean of the overall counties, and with a small employment rate. In contrast, in countries of the associative positive stratum, the percentage of Hispanics is lower than the average, in favor of a higher percentage of whites and Asians, while the dissociative stratum is characterized by a higher percentage of Hispanics and a smaller community of black people, with a higher level of education. Furthermore, the associative positive stratum is distinguished by a high smoke rate. In contrast, the associative negative stratum identifies urban counties with a high population density. This population is characterized by higher income, but also a higher percentage of poor people and lower levels of education. The meteorological variable, such as the averages of daily temperatures, the relative average of humidity, and the dew point, appears to play an important role in the distinction of strata—that is, in the characterization of the different effects of the implementation of environmental plans on the level of pm2.5. We refer to the Supplementary Materials for more details about the probability of each county to belong in each of the three strata.
6 Conclusion
In this paper, we proposed the Confounders-Aware SHared Atoms Bayesian Mixture Model (CASBAH). CASBAH uses the dependent Dirichlet process to create an highly flexible and adaptable distribution for the potential post-treatment variables, taking into account both the confounders and the treatment variable. This approach extends the recent BNP model for causal inference proposed by Zorzetto et al. (2024) to the principal stratification settings. CASBAH also provides an estimate of causal effects for a set of innovative estimates of principal causal effects.
The proposed model improves nonparametric modeling for principal stratification, e.g., the model proposed by Schwartz et al. (2011), by offering a more flexible approach to handling post-treatment variables. CASBAH does so by integrating the confounders into the weights of an infinite Bayesian mixture model, improving control over measured confounding, and enhancing the precision of causal effect estimates. Additionally, by utilizing shared random kernel atoms, the model facilitates the identification of associative or dissociative strata without predetermined criteria, effectively learning from the data across different treatment levels.
CASBAH dynamically identifies the main strata and imputes missing post-treatment variables and outcomes without the need to set predefined thresholds on the post-treatment values, as is done in most of the principal stratification literature in the context of continuous post-treatment. Moreover, the model quantifies the uncertainty in membership in the principal strata and estimates the principal causal effects. Theorem 1 provides a useful factorization of the expected value for each potential outcome, while Theorem 2 illustrates how the probability of a unit belonging to a dissociative stratum is non-zero and depends on its observed covariates.
The performance of the proposed model is evaluated in a battery of Monte Carlo simulations where the performance of CASBAH is compared and contrasted with the performance of previous BNP modeling for principal stratification proposed by Schwartz et al. (2011). Furthermore, the proposed model is used to assess the effectiveness of a previous revision of NAAQS in reducing levels of pm2.5and, in turn, reducing the mortality rate in treated counties. Specifically, we find significant effects on the reduction of mortality rate in areas where pollution was effectively reduced by the non-attainment designation (associative negative effects). In contrast, those areas where pollution was not measurably affected (dissociative effect) or where it actually increased (associative positive effects) did not show any significant effect on the mortality rate.
Future research could extend our methodology by adopting more flexible approaches to model the final outcome. Although CASBAH currently employs a linear model for this outcome, future work could move beyond linearity to embrace more sophisticated modeling techniques. In this contribution, we focus on innovations in modeling the post-treatment variable and decided to avoid overcomplicating the outcome model. However, exploring more complex options for outcome models remains an opportunity for further research. Furthermore, our methodology could be adapted to accommodate continuous treatments, similar to the approach suggested by Antonelli et al. (2023). We also leave the exploration of this extension for future studies.
SUPPLEMENTARY MATERIAL TO
“BAYESIAN NONPARAMETRICS FOR PRINCIPAL STRATIFICATION WITH CONTINUOUS POST-TREATMENT VARIABLES”
DAFNE ZORZETTO, ANTONIO CANALE, FABRIZIA MEALLI, FRANCESCA DOMINICI, AND FALCO J. BARGAGLI-STOFFI
Appendix A Conjugate prior distribution for logit regression parameters
[…]
The SUN density distribution for $\beta^{(t)}$ in Eq. (7) is defined, for $q=(p+1)(L-1)$, as
$\mathds{P}(\beta^{(t)})=\phi_{q}(\beta^{(t)}-\xi;\Omega)\frac{\Phi_{h}(\gamma+\Delta^{T}\bar{\Omega}^{-1}\omega^{-1}(\beta^{(t)}-\xi);\Gamma-\Delta^{T}\bar{\Omega}^{-1}\Delta)}{\Phi_{h}(\gamma;\Gamma)},$ (A.1)
[…]
arellano2006unification show that if $\beta^{(t)}\sim\textsc{SUN}_{q,h}(\xi,\Omega,\Delta,\gamma,\Gamma)$ then
$\beta^{(t)}$ =$\stackrel{{\scriptstyle d}}{{=}}\xi+\omega(B_{0}^{(t)}+\Delta\Gamma^{-1}B_{1}^{(t)}),$
$B_{0}^{(t)}$ $\sim\mathcal{N}_{q}(0,\bar{\Omega}-\Delta\Gamma^{-1}\Delta^{T}),$
$B_{1}^{(t)}$ $\sim TN_{h}(-\gamma;0,\Gamma),$
where $TN_{h}(-\gamma;0,\Gamma)$ denotes an $h$-variate Gaussian with zero mean, covariance matrix $\Gamma$ and truncation below $-\gamma$. A simple mechanism that helps in the simulation of SUN variables.
The multinomial probit distribution for the weights $\pi^{(t)}=\{\pi_{l}^{(t)}\}_{l=1}^{L}$ can be rewritten as
$\mathds{P}(S_{i}(t)=l|\beta^{(t)},X_{i})=\Phi(x_{i}^{T}\beta_{l}^{(t)})\prod_{k=1}^{l-1}[1-\Phi(x_{i}^{T}\beta_{k}^{(t)})]=\prod_{k=1}^{l}\Phi\left((2\bar{s}_{ik}^{(t)}-1)x_{i}^{T}\beta_{k}^{(t)}\right)=\Phi_{l}(x_{i}^{T}\beta^{(t)};\mathbf{I}_{l})$
for $t=\{0,1\}$ and $l=1,\dots,L-1$, and where $x_{i}=(1,x_{i1},\dots,x_{ip}^{T})$ is the vector of the $p$ covariates and intercept for the unit $i$, $\bar{s}_{i}^{(t)}=(0_{S_{i}(t)-1}^{T},1)^{T}$ if $S_{i}(t)\leq L-1$ and $\bar{s}_{i}^{(t)}=0_{L-1}$ if $S_{i}(t)=L$, and $\mathbf{I}_{l}$ refers to the $l\times l$ identity matrix.
Consequently, the probability over the observation $i=1,\dots,n_{t}$ for $t=\{0,1\}$ is
$\mathds{P}(S^{(t)}|\beta^{(t)},X)=\prod_{i=1}^{n^{(t)}}\mathds{P}(S^{(t)}|\beta^{(t)},X_{i})=\Phi_{\bar{n}_{t}}(\bar{X}^{(t)}\beta^{(t)},\mathbf{I}_{\bar{n}^{(t)}})$ (A.2)
where $\bar{n}^{(t)}=n_{1}^{(t)}+\dots+n_{n}^{(t)}$ with $n_{i}^{(t)}=\min(S_{i}(t),L-1)$, $\bar{X}^{(t)}$ is a $\bar{n}^{(t)}\times[(p+1)(L-1)]$ matrix with row blocks $\bar{X}^{(t)}_{[i]}=X_{i}^{(t)}$ and $X_{i}^{(t)}=(diag(2\bar{s}_{i}^{(t)}-1)\bigotimes x_{i}^{T},0_{(n_{i}^{(t)}\times[(p+1)(L-1-n_{i}^{(t)})])})$.
Considering with the prior (A.1) and the likelihood (A.2), the posterior distribution for $\beta^{(t)}$ is
$\mathds{P}(\beta^{(t)}|S^{(t)},X)=\phi_{q}(\beta^{(t)}-\xi;\Omega)\frac{\Phi_{h+{\bar{n}_{t}}}(\gamma_{pst}+\Delta_{pst}^{T}\bar{\Omega}^{-1}\omega^{-1}(\beta^{(t)}-\xi);\Gamma_{pst}-\Delta_{pst}^{T}\bar{\Omega}^{-1}\Delta_{pst})}{\Phi_{h+{\bar{n}_{t}}}(\gamma_{pst};\Gamma_{pst})},$ (A.3)
where $\Delta_{pst}=(\Delta,\bar{\Omega}\omega(\bar{X}^{(t)})^{T}d^{-1})$, $\gamma_{pst}=(\gamma^{T},\xi^{T}(\bar{X}^{(t)})^{T}d^{-1})$, $\Gamma_{pst}$ is an $(h+{\bar{n}_{t}})\times(h+{\bar{n}_{t}})$ covariance matrix with blocks $\Gamma_{pst[11]}=\Gamma$, $\Gamma_{pst[22]}=d^{-1}(\bar{X}^{(t)}\Omega(\bar{X}^{(t)})^{T}+\mathbf{I}_{\bar{n}^{(t)}})d^{-1}$, and $\Gamma_{pst[12]}=\Gamma_{pst[21]}=d^{-1}\bar{X}^{(t)}\omega\Delta$, where $d=[(\bar{X}^{(t)}\Omega(\bar{X}^{(t)})^{T}+\mathbf{I}_{\bar{n}^{(t)}})\bigodot\mathbf{I}_{\bar{n}^{(t)}}]^{1/2}$.
In the particular case in which the prior for $\beta^{(t)}$ is a multivariate Gaussian distribution—i.e.$h=0$—, then the posterior is still the SUN distribution in Eq. (A.3) with $\Delta_{pst}=\bar{\Omega}\omega(\bar{X}^{(t)})^{T}d^{-1}$, $\gamma_{pst}=d^{-1}\bar{X}^{(t)}\xi$, and $\Gamma_{pst}=d^{-1}(\bar{X}^{(t)}\Omega(\bar{X}^{(t)})^{T}+\mathbf{I}_{\bar{n}^{(t)}})d^{-1}$.
[…]
$\beta^{(t)}$ =$\stackrel{{\scriptstyle d}}{{=}}\xi+\omega(B_{0,pst}^{(t)}+\Delta_{pst}\Gamma_{pst}^{-1}B_{1,pst}^{(t)}),$
$B_{0,pst}^{(t)}$ $\sim\mathcal{N}_{q}(0,\mathbf{I}_{q}-\Delta_{pst}\Gamma_{pst}^{-1}\Delta_{pst}^{T}),$
$B_{1,pst}^{(t)}$ $\sim TN_{h+{\bar{n}^{(t)}}}(-\gamma_{pst};0,\Gamma_{pst}),$
with $\Delta_{pst}=\omega(\bar{X}^{(t)})^{T}d^{-1}$, $\gamma_{pst}=d^{-1}\bar{X}^{(t)}\xi$, and $\Gamma_{pst}=d^{-1}(\omega^{2}\bar{X}^{(t)}(\bar{X}^{(t)})^{T}+\mathbf{I}_{\bar{n}^{(t)}})d^{-1}$ where $d=[(\omega^{2}\bar{X}^{(t)}(\bar{X}^{(t)})^{T}+\mathbf{I}_{\bar{n}^{(t)}})\bigodot\mathbf{I}_{\bar{n}^{(t)}}]^{1/2}$.
Appendix B Proof Theorem 1
Theorem 1 in Section 3.1 wants to rewrite the causal estimands in (1) that are not identifiable.
[…]
$\mathbb{E}\big{[}Y_{i}(1)-Y_{i}(0)\mid S_{i}(1)=s_{1},S_{i}(0)=s_{0}\big{]}$
$\quad=\mathbb{E}\big{[}Y_{i}(1)\mid S_{i}(1)=s_{1},S_{i}(0)=s_{0}\big{]}-\mathbb{E}\big{[}Y_{i}(0)\mid S_{i}(1)=s_{1},S_{i}(0)=s_{0}\big{]}$
$\quad=\mathbb{E}_{X_{i}\mid S_{i}(1),S_{i}(0)}\mathbb{E}\big{[}Y_{i}(1)\mid S_{i}(1)=s_{1},S_{i}(0)=s_{0},X_{i}=x\big{]}$
$\quad\quad-\mathbb{E}_{X_{i}\mid S_{i}(1),S_{i}(0)}\mathbb{E}\big{[}Y_{i}(0)\mid S_{i}(1)=s_{1},S_{i}(0)=s_{0},X_{i}=x\big{]}$
$\quad=\mathbb{E}_{X_{i}\mid S_{i}(1),S_{i}(0)}\mathbb{E}\big{[}Y_{i}\mid T_{i}=1,S_{i}(1)=s_{1},S_{i}(0)=s_{0},X_{i}=x\big{]}$
$\quad\quad-\mathbb{E}_{X_{i}\mid S_{i}(1),S_{i}(0)}\mathbb{E}\big{[}Y_{i}\mid T_{i}=0,S_{i}(1)=s_{1},S_{i}(0)=s_{0},X_{i}=x\big{]},$
where the first and second equality invokes the properties of expectation and the third equality is due to the no unmeasured confounding and SUTVA assumption. The inner expectation of each integral, for the treatment value $t\in\{0,1\}$, can be decomposed as
$\mathbb{E}\big{[}Y_{i}\mid T_{i}=t,S_{i}(1)=s_{1},S_{i}(0)=s_{0},X_{i}=x\big{]}$
$\quad=\mathbb{E}_{P_{i}(1),P_{i}(0)\mid T_{i},S_{i}(1),S_{i}(0),X_{i}}\mathbb{E}\big{[}Y_{i}\mid T_{i}=t,S_{i}(1)=s_{1},S_{i}(0)=s_{0},X_{i}=x,P_{i}(1)=p_{1},P_{i}(0)=p_{0}\big{]}$
$\quad=\mathbb{E}_{P_{i}(1),P_{i}(0)\mid T_{i},S_{i}(1),S_{i}(0),X_{i}}\mathbb{E}\big{[}Y_{i}\mid T_{i}=t,X_{i}=x,P_{i}(1)=p_{1},P_{i}(0)=p_{0}\big{]},$
where the second equality invokes the property that the outcome $Y_{i}$ is independent to the latent cluster-allocation $(S_{i}(0),S_{i}(1))$ given the potential post-treatment variable $(P_{i}(0),P_{i}(1))$, by definition of the post-treatment variable model.
Therefore, we can rewritten it explicitly, as reported in Theorem 1,
$\mathbb{E}\big{[}Y_{i}\mid T_{i}=t,S_{i}(1)=s_{1},S_{i}(0)=s_{0},X_{i}=x\big{]}$
$\quad=\int_{p_{0}p_{1}}\mathbb{E}\big{[}Y_{i}\mid T_{i}=1,X_{i}=x,P_{i}(1)=p_{1},P_{i}(0)=p_{0}\big{]}$
×$\quad\quad\times Pr\big{(}P_{i}(1)=p_{1},P_{i}(0)=p_{0}\mid T_{i}=t,S_{i}(1)=s_{1},S_{i}(0)=s_{0},X_{i}=x\big{)}dp_{0}p_{1}dx$ (B.1)
The inner expectations $\mathbb{E}\big{[}Y_{i}\mid T_{i}=t,X_{i}=x,P_{i}(1)=p_{1},P_{i}(0)=p_{0}\big{]}$ is estimate with the outcome model $Y_{i}\mid T_{i},X_{i},P_{i}(1),P_{i}(0)$, as well the probability $\Pr\big{(}P_{i}(1)=p_{1},P_{i}(0)=p_{0}\mid T_{i}=t,S_{i}(1)=s_{1},S_{i}(0)=s_{0},X_{i}=x\big{)}$, that is defined by the model for the potential post-treatment variables.
Appendix C Proof Dissociative Stratum Probability
As defined in Section 3.1, a unit $i$ is allocated in the dissociative stratum when the two latent cluster allocation variables have same value, i.e., $S_{i}(0)=S_{i}(1)$. Given the prior distribution $G^{(0)}_{x_{i}},G^{(1)}_{x_{i}}$, the probability of unit $i$ to be allocated in the dissociative stratum is the following:
$\Pr[S_{i}(0)=S_{i}(1)]$ $=\mathbb{E}\big{[}\Pr[S_{i}(0)=S_{i}(1)\mid G^{(0)}_{x_{i}},G^{(1)}_{x_{i}}]\big{]}$
$=\mathbb{E}\bigg{[}\sum_{l=1}^{L}\pi_{l}^{(0)}(x_{i})\pi_{l}^{(1)}(x_{i})\bigg{]}$
$=\sum_{l=1}^{L}\mathbb{E}\big{[}\pi_{l}^{(0)}(x_{i})\pi_{l}^{(1)}(x_{i})\big{]}$
$=\sum_{l=1}^{L}\mathbb{E}\big{[}\pi_{l}(x_{i})^{2}\big{]},$ (C.1)
where the first and third equalities invokes the properties of expectation and the second equality invokes the definition of the latent variables $S_{i}^{(t)}$, for $t=\{0,1\}$, and the independence of the sequence of the random weights $\{\pi^{(0)}(x_{i})\}_{l\geq 1}$ and $\{\pi^{(1)}(x_{i})\}_{l\geq 1}$. Moreover the two sequence of the random weights are equivalent in distribution such that we can write $\pi_{l}^{(0)}(x)\stackrel{{\scriptstyle d}}{{=}}\pi_{l}^{(1)}(x)\stackrel{{\scriptstyle d}}{{=}}\pi_{l}(x)$, notation used in the fourth equality.
By the definition of the probit stick-breaking, the expected value of the $l$-th weight squared is the follows
$\mathbb{E}\big{[}\pi_{l}(x_{i})^{2}\big{]}$ $=\mathbb{E}\bigg{[}\Phi\big{(}\alpha_{l}(x_{i})\big{)}^{2}\prod_{r<l}\bigg{(}1-\Phi\big{(}\alpha_{r}(x_{i})\big{)}\bigg{)}^{2}\bigg{]}$
$=\mathbb{E}\bigg{[}\Phi\big{(}\alpha_{l}(x_{i})\big{)}^{2}\prod_{r<l}\bigg{(}1-\Phi\big{(}\alpha_{r}(x_{i}))^{2}-2\Phi\big{(}\alpha_{r}(x_{i})\big{)}\bigg{)}\bigg{]}$
$=\mathbb{E}\big{[}\Phi\big{(}\alpha_{l}(x_{i})\big{)}^{2}\big{]}\prod_{r<l}\bigg{(}1+\mathbb{E}\big{[}\Phi\big{(}\alpha_{r}(x_{i}))^{2}\big{]}-2\mathbb{E}\big{[}\Phi\big{(}\alpha_{r}(x_{i})\big{)}\big{]}\bigg{)}\bigg{]}$
For simplify the notation we will indicate $\rho_{1}(x_{i})=\mathbb{E}\big{[}\Phi\big{(}\alpha_{l}(x_{i})\big{)}\big{]}$ and $\rho_{2}(x_{i})=\mathbb{E}\big{[}\Phi\big{(}\alpha_{l}(x_{i})\big{)}^{2}\big{]}$, for each $l\in\{1,\dots,L\}$. Then (C.1) is rewritten as
$\Pr[S_{i}(0)=S_{i}(1)]$ $=\sum_{l=1}^{L}\mathbb{E}\big{[}\pi_{l}(x_{i})^{2}\big{]}$
$=\sum_{l=1}^{L}\bigg{[}\rho_{2}(x_{i})\prod_{r<l}\big{(}1+\rho_{2}(x_{i})-2\rho_{1}(x_{i})\big{)}\bigg{]}$
$=\sum_{l=1}^{L}\bigg{[}\rho_{2}(x_{i})(1+\rho_{2}(x_{i})-2\rho_{1}(x_{i}))^{l-1}\bigg{]}$
$=\frac{\rho_{2}(x_{i})[(1+\rho_{2}(x_{i})-2\rho_{1}(x_{i}))^{L}-1]}{\rho_{2}(x_{i})-2\rho_{1}(x_{i})},$
where the third and fourth equalities are following by properties of mathematical series, and it is reported as Theorem 2.
Appendix D Posterior Computation
[…]
In this section, we describe the Gibbs sampling algorithm for model fitting that allows us to draw from the posterior distribution. Following the steps in the algorithm 1, in each iteration $r=1,\dots,R$, we use the observed data $(y,p,t,x)$ to update the parameters and the augmented variables and impute the missing post-treatment variable $P^{mis}$ and missing outcome $Y^{mis}$.
The Gibbs sampling algorithm is divided into three parts: the estimation of the shared atoms mixture model for the post-treatment variables (divided in the estimation of cluster allocation, cluster-specific parameters, and confounder-dependent weights), the imputation of the missing post-treatment variables, and the estimation of the outcome model.
As already discussed, the outcome model is not our main concern, therefore we assume a linear model. In particular, in the following Gibbs sampler, we consider the outcome model used in the simulation study—i.e., (10)—where only the potential post-treatment variables are included. This choice is driven by the purpose of focusing attention on the essential definition of the relation between the post-treatment variable and outcome, which is crucial to impute the missing post-treatment variable $P^{mis}$. However, the algorithm can be easily modified to include the covariates $X$ in the linear regression or to consider a more complex and flexible model.
Cluster Allocation. The latent variables $S_{i}(t)$ identifies the cluster allocation for each units $i\in\{1,\dots,n\}$ at the treatment level $t$. Its posterior distribution is a multinomial distribution where
$\Pr\{S_{i}(t)=l\}\propto\pi_{l}^{(t)}({x}_{i})\mathcal{N}(p_{i};\eta_{l},\sigma_{l}^{2}),$
for $i=1,\dots,n$ and $l=1,\dots,L$, with $\pi^{(t)}_{l}$ defined as:
$\pi_{l}^{(t)}({x}_{i})=\Phi(\alpha_{l}^{(t)}({x}_{i}))\prod_{r<l}(1-\Phi(\alpha_{r}^{(t)}({x}_{i}))),$
for $l=1,\dots,L-1$ and with $\Phi(\alpha_{L}^{(t)}({x}_{i}))=1$.
Cluster Specific Parameters. Thanks to the latent variables $\{S_{i}(0),S_{i}(1)\}$, that cluster the units by the value of their outcome, we know for each cluster $l\in\{1,\dots,L\}$, the allocated units and we can update the values of the parameters from their posterior distributions:
$\eta_{l}$ $\sim\mathcal{N}\left(V_{l}^{-1}\times\left(\frac{\sum_{\{i:S_{i}(0)=l\}}p_{i}(0)+\sum_{\{i:S_{i}(1)=l\}}p_{i}(1)}{\sigma_{l}^{2}}+\frac{\mu_{\eta}}{\sigma^{2}_{\eta}}\right),V_{l}^{-1}\right);$
$\sigma_{l}^{2}$ $\sim\mbox{InvGamma}\left(\gamma_{1}+\frac{n_{l}}{2},\gamma_{2}+\frac{\sum_{\{i:S_{i}(0)=l\}}(p_{i}(0)-\eta_{l})^{2}+\sum_{\{i:S_{i}(1)=l\}}(p_{i}(1)-\eta_{l})^{2}}{2}\right);$
for $l=1,\dots,L$ and where $V_{l}=n_{l}/\sigma_{l}^{2}+1/\sigma^{2}_{\eta}$ and $n_{l}$ is the number of units allocated in the $l$-th cluster.
Confounder-Dependent Weights. The $\{\beta_{ql}^{(t)}\}_{q=0}^{p}=(\beta_{0l}^{(t)},\beta_{l}^{(t)})$, for $l=1,\dots,\max(S_{i}(t),L-1)$, are updated for the posterior distribution:
$\beta^{(t)}$ =$\stackrel{{\scriptstyle d}}{{=}}\xi+\omega(B_{0,pst}^{(t)}+\Delta_{pst}\Gamma_{pst}^{-1}B_{1,pst}^{(t)}),$
$B_{0,pst}^{(t)}$ $\sim\mathcal{N}_{q}(0,\mathbf{I}_{q}-\Delta_{pst}\Gamma_{pst}^{-1}\Delta_{pst}^{T}),$
$B_{1,pst}^{(t)}$ $\sim TN_{h+{\bar{n}^{(t)}}}(-\gamma_{pst};0,\Gamma_{pst}),$
with $\Delta_{pst}=\omega(\bar{X}^{(t)})^{T}d^{-1}$, $\gamma_{pst}=d^{-1}\bar{X}^{(t)}\xi$, and $\Gamma_{pst}=d^{-1}(\omega^{2}\bar{X}^{(t)}(\bar{X}^{(t)})^{T}+\mathbf{I}_{\bar{n}^{(t)}})d^{-1}$ where $d=[(\omega^{2}\bar{X}^{(t)}(\bar{X}^{(t)})^{T}+\mathbf{I}_{\bar{n}^{(t)}})\bigodot\mathbf{I}_{\bar{n}^{(t)}}]^{1/2}$. More details for $\bar{X}^{(t)}$ and $\omega$ definitions in Section A.
Imputation Missing Post-Treatment Variables. For each unit $i\in\{1,\dots,n\}$, we impute the missing post-treatment variable $P_{i}^{mis}$. Firstly, drawing the relative cluster-allocation variable $S_{i}(1-t)$—where $t$ is the observed treatment of the unit $i$—from a multinomial distribution with
$\mathbb{P}\{S_{i}(1-t)=l\}\propto\pi_{l}^{(1-t)}({x}_{i}),$
for $l=1,\dots,L$. Where $\pi_{l}^{(1-t)}({x}_{i})=\Phi(\alpha_{l}^{(1-t)}({x}_{i}))\prod_{r<l}(1-\Phi(\alpha_{r}^{(1-t)}({x}_{i})))$, for $l=1,\dots,L-1$ and with $\Phi(\alpha_{L}^{(1-t)}({x}_{i}))=1$.
Successively, drawing the missing post-treatment variable $P_{i}^{mis}$, conditioned to the allocation to the cluster $l$ and the observed outcome variables $Y_{i}(t)$. For each $i$ such that the observed treatment level is $T=1$, $P_{i}(1-t)$ is drown from
$\{P_{i}^{mis}|S_{i}(1-t)=l,\eta,\sigma^{2},P_{i},Y_{i}\}\sim\mathcal{N}\left(v^{-1}\left(\frac{\eta_{l}}{\sigma^{2}_{l}}+\frac{m_{1}}{v_{1}}\right),v^{-1}\right);$
where
$v=\frac{1}{\sigma^{2}_{l}}+\frac{1}{v_{1}},\quad m_{1}=\frac{Y_{i}(1)-\theta_{10}-\theta_{11}P_{i}(1)}{\theta_{12}+\theta_{13}P_{i}(1)},\quad v_{1}=\frac{e^{\lambda_{0}+\lambda_{1}P_{i}(1)}}{(\theta_{12}+\theta_{13}P_{i}(1))^{2}}.$
While for each $i$ such that the observed treatment level is $T=0$, $P_{i}(1-t)$ is drown from
$\{P_{i}^{mis}|S_{i}(1-t)=l,\eta,\sigma^{2},P_{i},Y_{i}\}\sim\mathcal{N}(\eta_{l},\sigma^{2}_{l}).$
Outcome Model. The $\theta^{(t)}$ parameters are independent for the treatment level $t$, therefor the posterior distributions are, respectively for $t=0,1$:
$\theta^{(t)}$ $\sim\mathcal{N}_{q^{(t)}}((V^{(t)})^{-1}M^{(t)},V^{(t)})^{-1});$
$M^{(t)}$ $=(\tilde{P}^{(t)})^{T}\Phi^{(t)}\tilde{P}^{(t)}+(\sigma^{2}_{\theta})^{-1}\mathbf{I}_{q^{(t)}};$
$V^{(t)}$ $=(\tilde{P}^{(t)})^{T}\Phi^{(t)}Y(t)+\frac{\mu_{\theta}}{\sigma^{2}_{\theta}}.$
For the treatment level $t=0$: $q^{(0)}=2$; $\tilde{P}^{(0)}$ is a matrix $n_{0}\times n_{0}$ such that $\tilde{P}^{(0)}=[1_{n_{0}},P(0)]$ with $1_{n_{0}}$ a vector of $1$ and $P(0)$ the vector of observed values of post-treatment variable for the units $n_{0}$ assigned at the control group—i.e. $t=0$; and $\Phi^{(0)}$ is a diagonal matrix $n_{0}\times n_{0}$ with value $\exp(\lambda_{0})$ in the diagonal. In similar way, for the treatment level $t=1$: $q^{(1)}=4$; $\tilde{P}^{(1)}$ is a matrix $n_{1}\times n_{1}$ such that $\tilde{P}^{(1)}=[1_{n_{0}},P(1),P(0),P(1)\cdot P(0)]$ with $1_{n_{1}}$ a vector of $1$, $P(1)$ the vector of observed values of post-treatment variable for the units $n_{1}$ assigned at the treated group—i.e. $t=0$—and $P(0)$ the vector of imputed values of post-treatment variable; and $\Phi^{(1)}$ is a diagonal matrix $n_{1}\times n_{1}$ with the values $\exp(\lambda_{0}+\lambda_{1}P(1))$ in the diagonal.
The parameters in the variance of the $Y$-model, $\lambda_{0}$ and $\lambda_{1}$, do not have conjugate priors, therefore a independent Metropolis proposal step is necessary. At each iteration $r\in\{1,\dots,R\}$, $\lambda_{0}^{*}$ and $\lambda_{1}^{*}$ are drown from the proposal distribution $\mathcal{N}(\mu_{\lambda_{0}},\sigma^{2}_{\lambda_{0}})$ and $\mathcal{N}(\mu_{\lambda_{1}},\sigma^{2}_{\lambda_{1}})$ respectively. Then, at iteration $r$ the value of the parameter are updated as following: $\lambda_{0}^{(r)}=\lambda_{0}^{*}$ with probability
$\frac{\prod_{i\in n}\mathcal{N}(Y_{i}|\mu_{Y}^{(t)},\exp(\lambda_{0}^{*}+\mathbb{I}_{(T_{i}=1)}\lambda_{1}^{(r-1)}P_{i}(1))}{\prod_{i\in n}\mathcal{N}(Y_{i}|\mu_{Y}^{(t)},\exp(\lambda_{0}^{(r-1)}+\mathbb{I}_{(T_{i}=1)}\lambda_{1}^{(r-1)}P_{i}(1))},$
otherwise $\lambda_{0}^{(r)}=\lambda_{0}^{(r-1)}$; and $\lambda_{1}^{(r)}=\lambda_{1}^{*}$ with probability
$\frac{\prod_{i\in n_{1}}\mathcal{N}(Y_{i}|\mu_{Y}^{(t)},\exp(\lambda_{0}^{(r-1)}+\lambda_{1}^{*}P_{i}(1))}{\prod_{i\in n_{1}}\mathcal{N}(Y_{i}|\mu_{Y}^{(t)},\exp(\lambda_{0}^{(r-1)}+\lambda_{1}^{(r-1)}P_{i}(1))},$
otherwise $\lambda_{1}^{(r)}=\lambda_{1}^{(r-1)}$; where $\mu_{Y}^{(0)}=\theta_{00}+\theta_{01}P_{i}(0)$ and $\mu_{Y}^{(1)}=\theta_{10}+\theta_{11}P_{i}(1)+\theta_{12}P_{i}(0)+\theta_{13}P_{i}(0)P_{i}(1)$.
Algorithm 1 Confounders-Aware Shared-atoms Bayesian Mixture Model
Inputs:
- the observed data $(y,p,t,x)$.
Outputs:
- posterior distributions of parameters: $\eta$, $\sigma$, $\beta$, $\theta$, and $\lambda$;
- imputed values for $P^{mis}$;
- posterior distribution over the space of partitions of the units.
Procedure:
Initialization of all parameters and latent variables.
For $r\in\{1,\dots,R\}$ :
$\xrightarrow{}$ Estimation of Shared Atoms Mixture Model:
Compute $\omega^{(t)}(x_{i})$ for $i=1,\dots,n$ and $t=0,1$;
Draw $S_{i}(t)$ for $i=1,\dots,n$ and $t=0,1$;
Draw $\eta$ and $\sigma$;
Compute $\alpha^{(t)}({x}_{i})$ for $i=1,\dots,n$ and $t=0,1$;
[…]
$\xrightarrow{}$ Imputation of Missing Post-Treatment Variables:
Draw $P_{i}^{mis}$ for $i=1,\dots,n$ and $t=0,1$.
$\xrightarrow{}$ Estimation of Outcome Model:
Draw $\theta^{(t)}$ for $t=0,1$;
Draw $\lambda_{0}$ and $\lambda_{1}$.
End
Appendix E More simulations details
Figure E.1 reports the distribution of the simulated post-treatment variables and outcomes for the five simulated scenarios. The complexity of the distributions increases over the simulated scenarios, in particular in scenarios 4 and 5, it is not unidentifiable the groups in the distributions of the potential post-treatment variables and the potential outcomes.
Figure E.1: Distributions of (right) the post-treatment variables and (left) the outcomes, for the five simulated scenarios. In light blue the distribution of the variables given the treatment and in yellow given the control.
Figure E.2 reports the results for the five simulated scenarios obtained with CASBAH. On the left, the boxplots show the distribution over the simulated samples of the expected values of the difference of the post-treatment variables under treatment and control in each stratum. The graphics confirm the ability of our proposed model to (i) identify correctly the number of strata—two strata in the simulated scenario 1 and 3, and three in the others—and (ii) capture the definition of the associative/dissociative strata without an a priori criteria—the dissociative stratum is always around zero for $\mathbb{E}[P(1)-P(0)]$, while the dissociative stratum does not include the zero. In the boxplots on the right, there is the distribution of the principal causal effect: EAE-, EDE, and EAE+. The different strata identify different treatment effects on the outcome, allowing us to characterize the heterogeneity in the causal effects. Few outliers are observed, however, they are found in particular in Scenario 5 which describes a more complex relation among variables and strata.
Figure E.2: Representations of the five simulated scenarios. (Left) The expected value of the difference of the post-treatment variables under treatment and control given the strata allocation. (Right) The expected associative/dissociative causal effects. In green and indicated with $V=-1$, the associative negative stratum—i.e. corresponding to $S_{i}(1)\prec S_{i}(0)$—, in yellow and indicated with $V=0$, the dissociative stratum—i.e., $S_{i}(1)=S_{i}(0)$—, and in red and indicated with $V=+1$ the associative positive stratum—i.e., $S_{i}(1)\succ S_{i}(0)$.
Appendix F More application details and results
Figure F.1 maps of the final analysis dataset described in Section 5.1 On the top map in Figure F.1, the red points visualize the treated counties, which appear to be closer to the main cities, such as Chicago, New York City, Washington DC or Cleveland. The continuous post-treatment variable is the difference in level pm2.5 between the follow-up period (2010-2016) and the baseline period (2000-2005), reported on the map to the left of the bottom of Figure F.1. The outcome variable is defined as the difference between the age-adjusted mortality rate between the follow-up period and the baseline period, and is visualized in the map at the bottom right of Figure F.1. As reported in the maps, both the post-treatment variable and the outcome have almost all negative values, highlighting a general trend to decrease the level of pm2.5 and a reduction in mortality rates in the last decade.
Figure F.1: Considered counties in the Easter U.S. (top) Attainment of the air pollution standard ($15\mu m$ for pm2.5) and consequent application of air pollution regulations: 00 if the county was in attainment, $1$ if the county was in non-attainment. (bottom left) Difference between the average long-term exposure pm2.5 between baseline and the follow-up period (measured in $\mu m$). (bottom right) Difference of the age-adjusted mortality rate between baseline and the follow-up period, value per 100,000.
In addition to the analysis reported in Section 5.2, our proposed approach allows us to quantify the uncertainty of strata allocation. In fact, for each county, we know the probability of being allocated in each of the three strata, in addition to the estimation of its allocation. In addition, we can visualize this information on the US map. Specifically, the first three maps in Figure F.2 visualize the probability that each county is assigned to the three different strata. As already underlined in Figure F.3, counties with a higher probability of being allocated in the associative negative stratum and the dissociative stratum are far from the largest cities, different from the associative positive stratum. In addition, western countries seem to have a small probability of being assigned to the associative negative stratum. The fourth map—bottom right in Figure F.2—reports the estimation of the partition point of the strata, a partition that is used to estimate the principal causal effects.
Figure F.2: Considered counties in the Easter U.S.. (top left) Probability to be allocated in the associative positive stratum. (top right) Probability to be allocated in the dissociative stratum. (bottom left) Probability to be allocated in the associative negative stratum. (bottom right) Point estimation of strata allocation.
Figure F.3: Representation of the characteristics of the identified strata. Each spider plot reports in the colored area the strata-specific characteristics—the mean of the analyzed covariates—and in the gray area the collective characteristics—the mean of the covariates among all the analyzed counties in the Eastern U.S.. We can consider the gray area as the benchmark to understand how the characteristics of each stratum differ from the collective characteristics of the analyzed population.
[…]