### v2

# AppReferences

Abstract
Principal stratification provides a causal inference framework for investigating treatment effects in the presence of a post-treatment variable. Principal strata play a key role in characterizing the treatment effect by identifying groups of units with the same or similar values for the potential post-treatment variable under both treatment levels. The literature has focused mainly on binary post-treatment variables. Few papers considered continuous post-treatment variables. In the presence of a continuous post-treatment, a challenge is how to identify and characterize meaningful coarsening of the latent principal strata that lead to interpretable principal causal effects. This paper introduces the confounders-aware shared-atom Bayesian mixture, a novel approach for principal stratification with binary treatment and continuous post-treatment variables. Our method leverages Bayesian nonparametric priors with an innovative hierarchical structure for the potential post-treatment variable that overcomes some of the limitations of previous works. Specifically, the novel features of our method allow for (i) identifying coarsened principal strata through a data-adaptive approach and (ii) providing a comprehensive quantification of the uncertainty surrounding stratum membership. Through Monte Carlo simulations, we show that the proposed methodology performs better than existing methods in characterizing the principal strata and estimating principal effects of the treatment. Finally, our proposed model is applied to a case study in which we estimate the causal effects of US national air quality regulations on pollution levels and health outcomes.
keywords:
Bayesian Causal Inference; Principal Stratification; Shared Atom Mixture Model.
1 Introduction
Principal stratification is a well-established causal inference approach that accounts for post-treatment variables (Frangakis & Rubin, 2002). A post-treatment variable is defined as a variable that is potentially affected by treatment and it is also associated with the primary potential outcomes. Thus, a post-treatment variable can be seen as an intermediate variable between the treatment and the outcome of interest. Post-treatment variables may pose inferential challenges for the estimation of treatment effects, but also provide a useful approach to conceptualizing the mechanisms/pathways through which the treatment affects the outcome through the intermediate variable.
Principal stratification with respect to the post-treatment variable, involves stratifying units into different groups, known as principal strata. Principal strata consist of units that share the same or similar potential values of the post-treatment variable under all possible treatment levels: by definition, principal strata are not affected by the treatment and thus represent an inherent latent characteristic of the units.
Since the assignment of treatment does not influence the principal strata, comparisons of potential outcomes at different treatment levels within a principal stratum represent well-defined causal effects (Rubin, 1974). These specific comparisons are referred to as principal causal effects. Researchers have often referred to the dissociative (and associative) effect as the effect of treatment on the outcome for the stratum that includes units whose potential post-treatment outcomes do not change (or change) under different treatment levels (Frangakis & Rubin, 2002; Mealli & Mattei, 2012).
Most of the literature on principal stratification and principal causal effects has focused on the case of binary treatment and the binary post-treatment variable (see, e.g., Angrist et al., 1996). This implies that researchers deal with at most four principal strata that are the possible combinations of values of these two variables (see the broad literature for binary or discrete post-treatment variables, e.g., Cheng & Small, 2006; Imai, 2008; Ding et al., 2011; Mattei & Mealli, 2011; Mealli & Pacini, 2013; Mealli et al., 2016; Jiang et al., 2016, 2022; Mattei et al., 2024). Considering continuous post-treatment variables leads to a potentially infinite number of principal strata, which, in turn, makes inference and interpretation very challenging. Various solutions have been proposed in the causal inference literature. Proposals include: (i) dichotomize the post-treatment variable (Sjölander et al., 2009; Jiang et al., 2022), or model the continuous post-treatment variable conditional on the binary treatment using (ii) fully parametric or semiparametric modeling (Jin & Rubin, 2008; Conlon et al., 2014; Lu et al., 2023), or (iii) nonparametric modeling (Schwartz et al., 2011).
Although these methods provide potential solutions to the challenges of continuous post-treatment variables, they often require strong assumptions or subjective choices. On the one hand, dichotomization of the post-treatment variable requires the choice of the threshold on the post-treatment variable (e.g., Zigler et al., 2012; Antonelli et al., 2023). Although this choice might be informed by the relevant subject matter literature, it entails a level of arbitrariness, and requires extensive checking of the robustness of the results with respect to the chosen thresholds.
Our proposed method introduces a novel and more interpretable approach to characterizing the principal strata, addressing some critical challenges in this framework. While, similarly Schwartz et al. (2011) modeled the continuous post-treatment variable using a Dirichlet process mixture (Ferguson, 1973), our approach advances this framework in two significant ways. First, we foster information sharing between the marginal distribution of potential post-treatment variables to enhance the ability to uncover the principal strata with a data-adaptive approach and provide a more interpretable definition of them. Second, while the Dirichlet process mixture proposed by Schwartz et al. (2011) adheres to the standard stick-breaking definition by Sethuraman (1994), our approach extends this by incorporating covariates into the weights of the mixture model. This adjustment significantly improves the imputation of missing potential post-treatment outcomes and enables strata identification without requiring predefined assumptions, consequently enhancing the estimation performance of treatment effects.
To do so, we introduce a novel Bayesian approach to principal stratification. Specifically, we introduce a confounders-aware shared-atoms Bayesian mixture model endowed with innovative estimands for principal causal effects. Our model exploits the dependent Dirichlet process (Quintana et al., 2022) to establish a highly flexible distribution for the potential post-treatment outcomes given the confounders and the treatment variable, similar to the recent proposal of Zorzetto et al. (2024) in the context of heterogeneity in the treatment effects.
Our proposed approach improves the current literature on nonparametric modeling for principal stratification in many respects. Firstly, it allows for the inclusion of the confounders in the weights of the infinite Bayesian mixture model, which provides a way to control for measured confounding and a more precise estimation of the causal effects. Secondly, the proposed model, which leverages shared random kernels’ atoms, allows us to identify associative/dissociative strata borrowing information across the treatment levels. Hence, our approach eliminates the need for subjective criteria to cluster principal strata. Instead, our model simultaneously (i) identifies the principal strata, according to the nature of the observed data, and (ii) imputes the missing post-treatment variables and outcomes. Moreover, our framework naturally incorporates uncertainty quantification for principal strata membership while providing robust estimation of principal causal effects.
This paper is organized as follows. In Section 2, we introduce the framework and the definition of the novel principal causal estimands. In Section 3, we propose our model and detail how it allows us to identify meaningful coarsening of the principal strata using a data-adaptive method. Theoretical prior properties provide insight into prior elicitation. An efficient Gibbs sampling algorithm to sample from the posterior distributions and estimate the model is also derived. Section 4 illustrates, via simulations, the ability of the proposed model to impute missing post-treatment outcomes, discover the main strata, and estimate the principal causal effects. In Section 5, we estimate the principal causal effects of air pollution policies on the mortality rate in the eastern United States (Zigler et al., 2018), where long-term fine particle exposure is the post-treatment variable. In Section 6, we conclude the paper with a discussion and avenues for future work.
2 Setup
We assume that we observe $n$ independent units. For each unit $i\in\{1,\dots,n\}$, let $\mathbf{X}_{i}\in\mathcal{X}\subseteq\mathbb{R}^{P}$ the set of observed covariates, $T_{i}\in\{0,1\}$ the observed (binary) treatment, $P_{i}\in\mathcal{P}\subseteq\mathbb{R}$ the post-treatment variable, and $Y_{i}\in\mathcal{Y}\subseteq\mathbb{R}$ the primary response. For each unit $i$, the potential post-treatment outcomes are defined as $\{P_{i}(0),P_{i}(1)\}\in{\mathds{R}}^{2}$ and the potential primary outcomes are defined as $\{Y_{i}(0),Y_{i}(1)\}\in{\mathds{R}}^{2}$, for $i=1,\dots,n$. The vector $\{P_{i}(0),P_{i}(1)\}$ represents the collection of the two potential values of the post-treatment variable, when the unit $i$ is assigned to the control or the treatment group, respectively. Similarly, the vector $\{Y_{i}(0),Y_{i}(1)\}$ represents the collection of the two potential values for the response variable, with $Y_{i}(0)$ when the unit $i$ is assigned to the control group and $Y_{i}(1)$ when assigned to the treatment group.
2.1 Identifying Assumptions
Following the principal stratification framework (Frangakis & Rubin, 2002; VanderWeele, 2011; Mealli & Mattei, 2012; Feller et al., 2017; Ding et al., 2017; Lu et al., 2023), the principal strata are not affected by treatment assignment; therefore, a principal stratification can be used as any unit classification to define meaningful causal estimates conditional on the principal strata and discover the heterogeneous treatment effect (Mealli & Mattei, 2012).
As highlighted in the Introduction, the definition of the dissociative and associative strata is of particular interest. Formally, a dissociative stratum includes units where the treatment does not affect the post-treatment variable, that is, $P_{i}(0)$ is the same as $P_{i}(1)$. In contrast, the associative stratum includes units for which the treatment affects the post-treatment variable, yielding a positive effect, positive associative stratum, or a negative effect, negative associative stratum.
Throughout the paper, we invoke the following assumptions:
{assumption}
[Stable Unit Treatment Value Assumption]
$Y_{i}(T_{1},T_{2},\cdots,T_{i},\cdots,T_{n})=Y_{i}(T_{i}),\quad Y_{i}(T_{i})=Y_{i},\mbox{ for }i=1,\dots,n;$
$P_{i}(T_{1},T_{2},\cdots,T_{i},\cdots,T_{n})=P_{i}(T_{i}),\quad P_{i}(T_{i})=P_{i},\mbox{ for }i=1,\dots,n.$
The Stable Unit Treatment Value Assumption (SUTVA) is a combination of (i) no interference between units, that is, the potential outcome and the potential values of the post-treatment variable of the unit $i$ do not depend on the treatment applied to other units, and (ii) consistency, that is, no different versions of the treatment levels assigned to each unit (Rubin, 1986). Under the principal stratification framework, this assumption is invoked for both the primary outcome variable and the post-treatment variable.
In practice, for $i=1,\dots,n$, we observe $p_{i}\in{\mathds{R}}$ and $y_{i}\in{\mathds{R}}$, that is, the realization of the random variables $P_{i}$ and $Y_{i}$, respectively, and defined as $P_{i}=(1-T_{i})\cdot P_{i}(0)+T_{i}\cdot P_{i}(1)$ and $Y_{i}=(1-T_{i})\cdot Y_{i}(0)+T_{i}\cdot Y_{i}(1)$.
{assumption}
[Strongly Ignorable Treatment Assignment] Given the observed covariates $\mathbf{x}_{i}$,
$\{Y_{i}(1),Y_{i}(0),P_{i}(0),P_{i}(1)\}\perp\!\!\!\perp T_{i}\mid X_{i},$
$0<Pr\left(T_{i}=1\mid X_{i}=x\right)<1\>\>\forall\>x\in\mathcal{X},$
where $\mathcal{X}$ is the features’ space. This assumption states that: (i) for each unit, the potential primary outcomes and the potential post-treatment variables are independent of the treatment conditional on the set of covariates $X_{i}$; (ii) all units have a positive chance of receiving the treatment.
3 Confounders-Aware Shared-atoms Bayesian Mixture Model
Following the Bayesian paradigm and invoking De Finetti’s theorem (Rubin, 1978), the joint probability distribution of the involved variables is unit exchangeable, and it is defined as
$\Pr(Y(0),Y(1),P(0),P(1),T,X)=\int_{\Theta}\prod_{i=1}^{n}\Pr(Y_{i}(0),Y_{i}(1),P_{i}(0),P_{i}(1),T_{i},X_{i},\theta)p(\theta)d\theta,$
where $p(\theta)$ is the prior distribution for all the involved parameters $\theta$ that take values in the parameter space $\Theta$. The inner probability can be factorized as
$\Pr(T_{i}\mid Y_{i}(0),Y_{i}(1),P_{i}(0),P_{i}(1),X_{i},\theta)$$\times\Pr(Y_{i}(0),Y_{i}(1)\mid P_{i}(0),P_{i}(1),X_{i},\theta)$
×$\times\Pr(P_{i}(0),P_{i}(1)\mid X_{i},\theta)$$\times\Pr(X_{i}\mid\theta).$ (1)
Invoking Assumption 2, the conditional probability of the treatment variable can be written as $\Pr(T_{i}\mid Y_{i}(0),Y_{i}(1),P_{i}(0),P_{i}(1),X_{i},\theta)=\Pr(T_{i}\mid X_{i},\theta)$. Moreover, we replace the conditional distribution of the covariates with the empirical distribution so that $\Pr(X_{i}\mid\theta)=\Pr(X_{i})$.
The remaining two probabilities in equation (1) must be modeled: (i) the distribution of potential primary outcomes conditional on the potential post-treatment variables and covariates, and (ii) the distribution of the potential post-treatment outcomes conditional on the covariates.
We focus our attention on the latter, that is, the distribution of the potential post-treatment outcomes conditional on the covariates for which we propose a novel Bayesian nonparametric approach. As our primary focus lies in modeling the post-treatment variable distribution, for the sake of clarity in our discussion, we employ a parametric model for the conditional response distribution following the settings of Schwartz et al. (2011). Note that, although the primary outcome model is assumed to be a linear regression model, it can also be generalized to a more complex and flexible model.
For modeling the post-treatment variable, we exploit a dependent nonparametric mixture, following the dependent Dirichlet process approach (MacEachern, 2000; Barrientos et al., 2012; Quintana et al., 2022; Denti et al., 2023). [^1] In particular, our model also shares some similarities with the hierarchical Dirichlet process (Teh et al., 2004, 2006; Teh & Jordan, 2009) and the recent common atom model of Denti et al. (2023). Specifically, we assume for each $i=1,\dots,n$:
$\{P_{i}\mid\mathbf{x}_{i},t\}$ $\sim f^{(t)}(\cdot\mid\mathbf{x}_{i}),\mbox{ for }t=\{0,1\},$
[…]
$\{G_{\mathbf{x}_{i}}^{(0)},G_{\mathbf{x}_{i}}^{(1)}\}$ $\sim\Pi,$ (2)
where ${\mathcal{K}}(\cdot;x,\psi)$ is a continuous density function, for every $\psi\in\Psi$, and $G_{\mathbf{x}_{i}}^{(t)}$ is a random probability measure depending on the confounders $x_{i}$ associated to an observation assigned to treatment level $t$. For each treatment level $t=\{0,1\}$, the distributions of $G_{\mathbf{x}_{i}}^{(0)}$ and $G_{\mathbf{x}_{i}}^{(1)}$ are characterized by specific weights assigned to a common set of atoms induced by the common process $\Pi$.
Following the characterization of the random probability measures, we write:
$G_{\mathbf{x}_{i}}^{(t)}=\sum_{l\geq 1}\pi_{l}^{(t)}(\mathbf{x}_{i})\delta_{\mathbb{\psi}_{l}},$ (3)
where the sequences $\{\pi_{l}^{(t)}(x_{i})\}_{l\geq 1}$ for $t=\{0,1\}$ represent infinite sequences of random weights, and $\{\psi_{l}\}_{l\geq 1}$ is an infinite sequence of random kernel’s parameters, independent and identically distributed from a base measure $H$, shared among potential post-treatment outcomes under both treatment levels $t$.
The sequences of weights are defined through a stick-breaking representation (Sethuraman, 1994),
$\pi_{l}^{(t)}(\mathbf{x}_{i})$ $=\omega_{l}^{(t)}(\mathbf{x}_{i})\prod_{r<l}\{1-\omega_{r}^{(t)}(\mathbf{x}_{i})\};$ (4)
where $\{\omega_{l}^{(t)}(\mathbf{x})\}_{l\geq 1}$ , for $t=\{0,1\}$ are $[0,1]$-valued independent stochastic processes.
The discrete nature of the random probability measure $G_{\mathbf{x}_{i}}^{(t)}$, for $t=\{0,1\}$, allows us to introduce the latent categorical variables $S_{i}^{(t)}$, that describe the cluster allocations for each unit $i\in\{1,\dots,n\}$, whose clusters are defined by heterogeneous values for $P_{i}(t)$. Assuming $\mathds{P}\{S_{i}^{(t)}=l\}=\pi_{l}^{(t)}(\mathbf{x}_{i})$, we can write model in (2), exploiting conditioning on $S_{i}^{(t)}$, as
$\{P_{i}|\mathbf{x}_{i},t,\psi,S_{i}^{(t)}=l\}\sim{\mathcal{K}}(\cdot\mid\mathbf{x}_{i},\psi_{l}),\quad\psi_{l}\sim H.$
where $\psi$ represents the infinite sequence $\{\psi_{l}\}_{l\geq 1}$.
Among the plethora of dependent non-parametric processes (see the recent review by Quintana et al., 2022, for a detailed description), we focused on the dependent probit stick-breaking process (Rodriguez & Dunson, 2011). Consistently with this, we specify:
$\omega_{l}^{(t)}(x_{i})=\Phi(\alpha_{l}^{(t)}(x_{i})),\quad\alpha_{l}^{(t)}(x_{i})\sim\mathcal{N}(\beta_{l0}^{(t)}+x_{i}^{T}\beta_{l}^{(t)},1),$ (5)
where $\Phi(\cdot)$ is the CDF of a standard Normal distribution, $\{\alpha_{l}^{(t)}(x_{i})\}_{l\geq 1}$ has Gaussian distributions with mean a linear combination of the $p$ covariates $x_{i}$, and $\{\beta_{l0}^{(t)},\beta_{l}^{(t)}\}_{l\geq 1}$ are the probit regression parameters treatment-specific.
We assume the kernel $\mathcal{K}$ to be Gaussian, so that model (2)–(3) becomes
$\{P_{i}(t)\mid\mathbf{x}_{i},S_{i}^{(t)}=l,\eta,\sigma\}\sim\mathcal{N}(\eta_{l},\sigma_{l}^{2}).$ (6)
where $\eta$ and $\sigma$ represent infinite sequences of location parameters $\{\eta_{l}\}_{l\geq 1}$ and scale parameters $\{\sigma_{l}\}_{l\geq 1}$, respectively, such that $\psi_{l}=(\eta_{l},\sigma_{l})$.
Prior elicitation is completed by assuming for the parameters $\eta_{l}$ and $\sigma_{l}$ in (6)
$\eta_{l}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(\mu_{\eta},\sigma_{\eta}^{2}),\mbox{ and }\sigma_{l}^{2}\stackrel{{\scriptstyle iid}}{{\sim}}\mbox{InvGamma}(\gamma_{1},\gamma_{2}).$
where InvGamma($\gamma_{1},\gamma_{2}$) represents the inverse gamma distribution with the shape parameter $\gamma_{1}\in\mathbb{R}^{+}$ and the scale parameter $\gamma_{2}\in\mathbb{R}^{+}$, and mean equal to $\gamma_{2}/(\gamma_{1}-1)$ and variance $\gamma_{2}^{2}/\{(\gamma_{1}-1)^{2}(\gamma_{1}-2)\}$.
For the regression parameters in (5), we assume the following multivariate Gaussian prior:
$\beta^{(t)}\sim\mathcal{N}_{(p+1)(L-1)}(\xi,\Omega)$ (7)
for $t=0,1$ and $l\geq 1$. According to Fasano et al. (2022), the Gaussian prior leads to a straightforward posterior computation, as discussed in the next section.
3.1 Causal Estimands
To define our causal estimands of interest, we postulate that the units can be marginally grouped into clusters with finite first moments with respect to the treatment level. Consequently, each unit $i$ can belong to: (i) the dissociative stratum if $S_{i}^{(1)}=S_{i}^{(0)}$, which is equivalent to $\mathbb{E}_{S_{i}^{(1)}}[P_{i}(1)]=\mathbb{E}_{S_{i}^{(0)}}[P_{i}(0)]$; or (ii) the associative positive stratum if $S_{i}^{(1)}\not=S_{i}^{(0)}$ and $\mathbb{E}_{S_{i}^{(1)}}[P_{i}(1)]>\mathbb{E}_{S_{i}^{(0)}}[P_{i}(0)]$ (throughout the paper, we indicate this condition as $S_{i}^{(1)}\succ S_{i}^{(0)}$); or (iii) the associative negative stratum if $S_{i}^{(1)}\not=S_{i}^{(0)}$ and $\mathbb{E}_{S_{i}^{(1)}}[P_{i}(1)]<\mathbb{E}_{S_{i}^{(0)}}[P_{i}(0)]$ (throughout, we indicate this condition as $S_{i}^{(1)}\prec S_{i}^{(0)}$).
Definition 3.1 (Causal Estimands).
Conditionally on the strata allocation, we define the causal estimands for the expected dissociative effect (EDE) and the expected associative effects (EAE) as
$EDE=\tau_{0}=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid S_{i}^{(1)}=S_{i}^{(0)}],$
$EAE_{+}=\tau_{+1}=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid S_{i}^{(1)}\succ S_{i}^{(0)}],$
$EAE_{-}=\tau_{-1}=\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid S_{i}^{(1)}\prec S_{i}^{(0)}].$
These causal estimands are not directly identifiable from the data due to the conditioning on unobservable quantities (i.e., $S_{i}^{(1)},S_{i}^{(0)}$). However, they can be weakly identified by invoking Assumptions (1) and (2).
Theorem 3.2.
Invoking assumptions (2.1) and (2.1), the expected value of each potential outcome can be rewritten as the following
$\mathbb{E}\big{[}Y_{i}(t)\mid S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big{]}$
$\quad=\int_{p_{0}p_{1}}\mathbb{E}\big{[}Y_{i}\mid T_{i}=t,X_{i}=x,P_{i}(1)=p_{1},P_{i}(0)=p_{0}\big{]}$
×$\quad\quad\times Pr\big{(}P_{i}(1)=p_{1},P_{i}(0)=p_{0}\mid T_{i}=t,S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big{)}dp_{0}p_{1}dx;$ (8)
where the inner expectations $\mathbb{E}\big{[}Y_{i}\mid T_{i}=t,X_{i}=x,P_{i}(1)=p_{1},P_{i}(0)=p_{0}\big{]}$ is estimate with the outcome model $Y_{i}\mid T_{i},X_{i},P_{i}(1),P_{i}(0)$, as well the probability $\Pr\big{(}P_{i}(1)=p_{1},P_{i}(0)=p_{0}\mid T_{i}=t,S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big{)}$, that is defined by the model for the potential post-treatment variables.
The proof is reported in the Supplementary Material.
3.2 Discovery of Principal Strata
While Bayesian nonparametric mixtures excel at capturing cluster structures, our focus shifts to leveraging this framework for causal inference. To estimate our causal estimands of interest, we need to characterize the joint distribution of $\{S_{i}^{(0)},S_{i}^{(1)}\}$ so that we can determine each unit’s allocation to a principal strata. This joint distribution cannot be directly observed due to the fundamental problem of causal inference. Nevertheless, our proposed Bayesian nonparametric model allows for robust estimation of this joint distribution through its hierarchical construction and shared parameters, effectively addressing this problem. In this section, we illustrate how our proposed Bayesian nonparametric model enables the characterization of principal strata in a statistically robust and fully data-adaptive manner.
We start from noting that the latent categorical variables $\{S_{i}^{(0)},S_{i}^{(1)}\}$ for each unit $i\in\{1,\ldots,n\}$ serve a dual purpose. First, they determine the allocation probabilities to specific mixture components. Second, they simultaneously define the marginal distributions of the random unit partition based on potential post-treatment outcomes (Quintana, 2006). This formulation creates a natural bridge between mixture model clustering and principal stratification, allowing us to characterize treatment effect heterogeneity through a principled probabilistic structure.
Our proposed model is defined such that the atoms $\{\psi_{l}\}_{l\geq 1}$ are shared between the two potential post-treatment outcomes for the same unit. This formulation yields precise posterior probabilities that quantify, for each unit $i$, whether the latent indicators $\{S_{i}^{(0)},S_{i}^{(1)}\}$ map to identical or distinct mixture components. Specifically, for any unit $i\in\{1,\ldots,n\}$, the probability of membership in the dissociative stratum—where $S_{i}^{(0)}\neq S_{i}^{(1)}$—equals the probability that the atom under control differs from the atom under treatment. This probability is provably non-zero and can be derived a priori through the following theorem.
Theorem 3.3.
Given a unit $i\in\{1,\dots,n\}$, with covariates $x_{i}$ following (2), the probability to belong to the dissociative stratum for each unit $i$ is a function of its covariates $x_{i}$, defined as:
$\Pr[S_{i}^{(0)}=S_{i}^{(1)}]=\frac{\rho_{2}(x_{i})[(1+\rho_{2}(x_{i})-2\rho_{1}(x_{i}))^{L}-1]}{\rho_{2}(x_{i})-2\rho_{1}(x_{i})}$ (9)
where $\rho_{1}(x_{i})=\mathbb{E}\big{[}\Phi\big{(}\alpha(x_{i})\big{)}\big{]}$ and $\rho_{2}(x_{i})=\mathbb{E}\big{[}\Phi\big{(}\alpha(x_{i})\big{)}^{2}\big{]}$ and $0\leq\Pr[S_{i}^{(0)}=S_{i}^{(1)}]\leq 1$.
The proof is reported in the Supplementary Material. The representation of the probability to belong to the dissociative stratum in (9) is further illustrated in Figure 1.
[…]
Figure 1: Probability to belong to the dissociative stratum. Function varies according to different values of $\alpha(x)=\beta_{0}+\beta_{1}x$.
3.3 Posterior Inference
Theorem 3.2 establishes that principal causal estimands are characterized as functions of both the primary response and potential post-treatment outcomes, each involving specific model parameters. Following the Bayesian paradigm, we estimate these parameters through their posterior distributions conditional on the observed data. This approach provides not only point estimates but also complete uncertainty quantification.
Sampling from the posterior joint distribution is straightforward via Gibbs sampling. In particular, the proposed algorithm (described in detail in the Supplementary Material) takes inspiration from Teh et al. (2004, 2006); Teh & Jordan (2009) for the hierarchical structure, and from Fasano et al. (2022) for the probit regression in the weights. We assume a general linear regression for the outcome model that depends on the parameter $\theta_{0}^{Y}$ for the outcome under control and $\theta_{1}^{Y}$ for the treated outcome.
The implementation of the Gibbs sampler proceeds through the following steps.
  • The latent variable $S_{i}^{(t)}$ is a multivariate distribution with
    $\Pr\{S_{i}^{(t)}=l\}\propto\pi_{l}^{(t)}(\mathbf{x}_{i})\Pr(P_{i}(t)\mid S_{i}^{(t)}=l,X_{i};\eta_{l},\sigma_{l}),$
    where $\pi_{l}^{(t)}(\mathbf{x}_{i})=\Phi(\alpha_{l}^{(t)}(\mathbf{x}_{i}))\prod_{r<l}(1-\Phi(\alpha_{r}^{(t)}(\mathbf{x}_{i})))p$ for $l=1,\dots,L-1$ and $\Phi(\alpha_{L}^{(t)}(\mathbf{x}_{i}))=1$.
  • The cluster-specific parameters are drawn from
    $\psi_{l}\propto\Pi_{x}\prod_{i:S_{i}^{(0)}=l}\Pr(P_{i}(0)\mid S_{i}^{(0)}=l,X_{i};\psi_{l})\prod_{i:S_{i}^{(1)}=l}\Pr(P_{i}(1)\mid S_{i}^{(1)}=l,X_{i};\psi_{l}).$
  • The logit regression parameters for conditional-dependent weights are drawn from
    $\beta^{(t)}\propto\Pr(\beta^{(t)})\prod_{i=1}^{n}\Pr(P_{i}(t)\mid S_{i}^{(t)},X_{i};\beta^{(t)})$
    following (Fasano et al., 2022) demonstration, detail in the Supplementary Material.
  • The missing post-treatment variable is imputed from
    $P_{i}(1-T_{i})\propto\Pr(Y_{i}\mid P_{i}(0),P_{i}(1),X_{i};\theta_{0}^{Y})^{1-T_{i}}\Pr(Y_{i}\mid P_{i}(0),P_{i}(1),X_{i};\theta_{1}^{Y})^{T_{i}}\Pr(P_{i}(1-T_{i})\mid X_{i};\psi).$
  • The posterior distribution for the outcome model is built from
    $\theta_{t}^{Y}\sim\Pr(\theta_{t}^{Y})\prod_{i=1}^{n}\text{P}(Y_{i}(t)\mid P_{i}(0),P_{i}(1),X_{i};\theta_{t}^{Y}).$
4 Simulation Study
The performance of the proposed mixture model is evaluated through a simulation study. Our objective is to investigate the model’s ability to (i) accurately impute the missing post-treatment and outcome variables, that is, to assess the potential bias of the expected value $E[P_{i}(1)-P_{i}(0)]$ and $E[Y_{i}(1)-Y_{i}(0)]$ over the units $i=1\dots,n$, (ii) correctly identify the principal strata and estimate the principal causal effects. To achieve this, we conducted simulations under five different data-generating mechanisms and analyzed the results to understand the model’s behavior in different scenarios.
The performance of our proposed approach is compared to those obtained with the Schwartz et al. (2011)’s model and the copula model proposed by Lu et al. (2023).
Commonly across the scenarios, we assume a linear regression model for the outcome model, defined as follows:
$\begin{bmatrix}Y(0)\\ Y(1)\end{bmatrix}\sim\mathcal{N}_{2}\begin{pmatrix}\begin{bmatrix}\theta_{00}+\theta_{01}P(0)\\ \theta_{10}+\theta_{11}P(1)+\theta_{12}P(0)+\theta_{13}P(0)P(1)\end{bmatrix},\begin{bmatrix}e^{\lambda_{0}}&0\\ 0&e^{\lambda_{0}+\lambda_{1}P(1)}\end{bmatrix}\end{pmatrix}.$ (10)
Assuming as prior distribution for the parameters
[…]
with $\mu_{\theta}=0$, $\sigma_{\theta}=10$, $\mu_{\lambda}=0$, and $\sigma_{\lambda}=2$. Clearly, our model can accommodate more complex choices for the outcome, but we decided to opt for a simpler model following what had already been done in the literature by Schwartz et al. (2011) and Lu et al. (2023).
For the data generating process, we simulate two Bernoulli confounders $(X_{1},X_{2})$ for scenarios 1 to 4 and five Bernoulli confounders $(X_{1},X_{2},X_{3},X_{4},X_{5})$ for scenario 5, and a binary treatment variable such that $T_{i}\sim\mbox{Be}(\mbox{logit}(0.4X_{1i}+0.6X_{2i}))$ for the scenarios $1-4$ and $T_{i}\sim\mbox{Be}(\mbox{logit}(0.4X_{1i}+0.6X_{2i}-0.3X_{3i}+0.2X_{4i}X_{5i}))$ for scenario 5, for $i=1,\dots,n$.
Each scenario assumes a different conformation of the strata for the continuous post-treatment variable $P_{i}=(P_{i}(0),P_{i}(1))\in\mathbb{R}^{2}$. Each stratum is obtained by introducing, for each unit, categorical variables $S_{i}^{(0)}$ and $S_{i}^{(1)}$, for control and treatment levels, respectively, with the vector of probabilities that depend on the values of the confounders and allocates the unit in different clusters. Conditionally on the cluster allocation $S_{i}^{(t)}=s$, with $s$ that has the same support for control and treatment levels, we simulate both potential post-treatment variables, under control and under treatment, respectively, as
$P_{i}(0)|S_{i}^{(0)}=s\sim\mathcal{N}(\eta_{s},\sigma^{2}_{s}),\quad P_{i}(1)|S_{i}^{(1)}=s\sim\mathcal{N}(\eta_{s},\sigma^{2}_{s}).$
Continuous potential outcomes $(Y_{i}(0),Y_{i}(1))$, for $i=1,\dots,n$, are simulated following the model (10) with common values for $\lambda_{0}=-0.5$ and $\lambda_{1}=0.1$, while $\theta^{(0)}$ and $\theta^{(1)}$ are different for each scenario. In each setting, the sample size is set to $n=500$. For each scenario, we simulate $100$ samples. Below we provide further details for each one of the five scenarios.
Scenario 1: We investigate a situation in which there are two strata: one with a dissociative effect and one with a positive associative effect. In particular, for $S_{i}^{(0)}=S_{i}^{(1)}=1$ $P_{i}(0),P_{i}(1)\sim\mathcal{N}(1,0.05)$, and for $S_{i}^{(0)}=2$ and $S_{i}^{(1)}=3$ $P_{i}(0)\sim\mathcal{N}(2,0.05)$ and $P_{i}(1)\sim\mathcal{N}(3,0.05)$. The regression parameters for the Y-model are $\theta^{(0)}=(1,2)$ and $\theta^{(1)}=(1,2,-1,0.5)$.
Scenario 2: We focus on a case where we have a dissociative stratum, for $S_{i}^{(0)}=S_{i}^{(1)}=1$ both $P(0)$ and $P(1)$ are simulated from $\mathcal{N}(2,0.05)$, an associative stratum with a positive effect, for $S_{i}^{(0)}=2$ and $S_{i}^{(1)}=3$ $P_{i}(0)\sim\mathcal{N}(2,0.05)$ and $P_{i}(1)\sim\mathcal{N}(3,0.05)$, and dissociative stratum with a negative effect, for $S_{i}^{(0)}=2$ and $S_{i}^{(1)}=1$ $P_{i}(0)\sim\mathcal{N}(2,0.05)$ and $P_{i}(1)\sim\mathcal{N}(1,0.05)$. The regression parameters for Y-model are $\theta^{(0)}=(1,2)$ and $\theta^{(1)}=(1,1.2,-1,1)$.
Scenario 3: This scenario corresponds to Scenario 1 with closer atoms for the strata and different variances. In particular, the dissociative stratum has $(P_{i}(0)|S_{i}^{(0)}=1)=(P_{i}(1)|S_{i}^{(1)}=1)\sim\mathcal{N}(1.5,0.12)$, and the associative stratum has $(P_{i}(0)|S_{i}^{(0)}=2)\sim\mathcal{N}(2,0.1)$ and $(P_{i}(1)|S_{i}^{(1)}=3)\sim\mathcal{N}(2.5,0.08)$. The regression parameters for Y-model are $\theta^{(0)}=(1,2)$ and $\theta^{(1)}=(1,1.2,-0.8,0.5)$.
Scenario 4: This scenario corresponds to Scenario 2 with closer atoms for the strata and different variances. In particular, the dissociative stratum has $(P_{i}(0)|S_{i}^{(0)}=1)=(P_{i}(1)|S_{i}^{(1)}=1)\sim\mathcal{N}(1.5,0.12)$, the associative positive stratum has $(P_{i}(0)|S_{i}^{(0)}=2)\sim\mathcal{N}(2,0.1)$ and $(P_{i}(1)|S_{i}^{(1)}=3)\sim\mathcal{N}(2.5,0.08)$, and the associative negative stratum has $(P_{i}(0)|S_{i}^{(0)}=2)\sim\mathcal{N}(2,0.1)$ and $(P_{i}(1)|S_{i}^{(1)}=1)\sim\mathcal{N}(1.5,0.12)$. The regression parameters for Y-model are $\theta^{(0)}=(1,2)$ and $\theta^{(1)}=(1,1.2,-0.8,0.5)$.
Scenario 5: We investigate the scenario with the three strata when the number of confounders increases, in particular the treatment variable and cluster allocation variables that depend on five confounders. The dissociative stratum has $(P_{i}(0)|S_{i}^{(0)}=1)=(P_{i}(1)|S_{i}^{(1)}=1)\sim\mathcal{N}(2,0.05)$, the associative positive stratum has $(P_{i}(0)|S_{i}^{(0)}=3)\sim\mathcal{N}(3,0.05)$ and $(P_{i}(1)|S_{i}^{(1)}=4)\sim\mathcal{N}(4,0.05)$, and the associative negative stratum has $(P_{i}(0)|S_{i}^{(0)}=2)\sim\mathcal{N}(2,0.05)$ and $(P_{i}(1)|S_{i}^{(1)}=1)\sim\mathcal{N}(1,0.05)$. The regression parameters for Y-model are $\theta^{(0)}=(1,2)$ and $\theta^{(1)}=(1,1.2,-1,0.5)$.
We choose the same hyperparameters for each setting such that the prior is non-informative and in common for all the settings. For the regression parameters in (5) and for the parameters $\eta_{l}$ and $\sigma_{l}$ in (6) we use the following conjugate priors
$\beta^{(t)}\sim\mathcal{N}_{(p+1)(L-1)}(0,20\times\mathbf{1}_{(p+1)(L-1)}),$
$\eta_{l}\sim\mathcal{N}(0,20),\mbox{ and }\sigma_{l}\sim\mbox{InvGamma}(2,0.5),$
for $t\in\{0,1\}$, $l\in\{1,\dots,20\}$, and $p$ according with the covariates considered in different settings, and where $\mathbf{1}_{q}$ is a diagonal matrix $q\times q$.
Table 4 reports the median and interquartile range of the bias for the expected value of the posterior distribution of the sample average of $P_{i}(1)-P_{i}(0)$ and $Y_{i}(1)-Y_{i}(0)$, for $i\in\{1,\dots,n\}$.
The results for the five scenarios demonstrate the strong ability of our proposed model to impute missing variables and accurately capture the true distribution across different data generating processes. The medians of the bias are close to zero and the interquartile range is reasonable given the simulated variability. In contrast, the two competing models highlight their limitations in recovering the same information. The algorithm from Lu et al. (2023) does not estimate the distribution of the potential outcome for the post-treatment variable, while the model from Schwartz et al. (2011) exhibits significant bias in four out of five scenarios and a higher interquartile range than our model across all scenarios. Furthermore, the comparison of $\mathbb{E}[Y_{i}(1)-Y_{i}(0)]$ reveals substantial bias in all five scenarios for both competing models Lu et al. (2023) and Schwartz et al. (2011).
[…]
Bias comparison of the two methods based on different simulation scenarios. Scenario 1 Scenario 2 Scenario 3 Scenario 4 Scenario 5 $\mathbb{E}[P(1)-P(0)]$ Our proposed method Median -0.0164 0.0027 -0.0318 0.0027 0.0011 IQR 0.0315 0.0254 0.0462 0.0325 0.0183 Schwartz et al. (2011) Median -0.0637 0.3388 0.1254 0.3280 0.3572 IQR 0.0989 0.0986 0.0956 0.0963 0.1216 Lu et al. (2024) Median IQR $\mathbb{E}[Y(1)-Y(0)]$ Our proposed method Median 0.0024 0.0256 0.0067 0.0419 0.0344 IQR 0.2235 0.1550 0.4661 0.1357 0.1316 Schwartz et al. (2011) Median -1.9891 -1.8265 -1.4363 -1.3115 -1.8856 IQR 0.2546 0.2110 0.1904 0.1836 0.3727 Lu et al. (2024) Median -1.3491 0.3070 -0.5848 0.1393 0.0751 IQR 0.1092 0.1206 0.1009 0.0869 0.1078 {tabnote} IQR: interquartile range.
To assess the accurate identification of the principal strata, we use the adjusted Rand index (Hubert & Arabie, 1985) and evaluate the estimation of the principal causal effects. The adjusted Rand index values for the five scenarios are reported in Table 4. For all scenarios, the index is close to $1$ and confirms that our proposed model can correctly identify the principal strata and, combined with the good missing data imputation, allows us to estimate the expected associative and dissociative effects. Table 4 shows that our model correctly identifies the number of principal strata—two in Scenario 1 and 3, three in the others—while also providing accurate causal effect estimates. In contrast, neither the model by Lu et al. (2023) nor that of Schwartz et al. (2011)identify the principal strata according to our definition.
[…]
Adjusted rand index for the five simulated scenarios computed on the point estimated partitions obtained with our proposed model. Scenario 1 Scenario 2 Scenario 3 Scenario 4 Scenario 5 Mean 0.9850 0.9906 0.9706 0.9717 0.9154 Standard deviation 0.0997 0.0597 0.1021 0.0892 0.1577
[…]
True and estimated principal causal effects—mean and standard deviation in the brackets—from our proposed method across scenarios. Scenario 1 Scenario 2 Scenario 3 Scenario 4 Scenario 5 True Estimated True Estimated True Estimated True Estimated True Estimated $\tau_{-}$ NA $-1.00$ $-0.98$ (0.32) NA $1.50$ $1.46$ (0.33) $-1.00$ $-0.79$ (0.68) $\tau_{0}$ $0.50$ $0.11$ (0.56) $4.00$ $4.03$ (0.61) $5.60$ $4.98$ (0.52) $4.00$ $4.11$ (0.53) $4.00$ $4.57$ (1.50) $\tau_{+}$ $9.00$ $8.79$ (1.25) $9.00$ $9.02$ (0.45) $8.60$ $9.36$ (0.87) $6.50$ $6.59$ (0.35) $9.00$ $8.91$ (0.44) {tabnote} NA: Not applicable. The stratum was not present in the scenario.
5 Assessing the Impact of Environmental Policies on Particulate Pollution and Health
As recently as 7 February 2024 the US Environmental Protection Agency (EPA) has proposed a revision of the National Ambient Air Quality Standards (NAAQS) (U.S. Environmental Protection Agency, 2024), which will result in a more stringent environmental policy for the concentration of air pollutants in the US that cause health hazards. To gather information and better understand the implementation of this revised policy, it is critically important to evaluate how past air quality regulations benefitted health outcomes and how these effects varied between different communities.
In our application, we explore this relevant research question by investigating the repercussions of a previous revision of NAAQS, leveraging insights gleaned from historical data and health data. First, we evaluate the direct and indirect effects of the 2005 NAAQS revision on the mortality rate in the period 2010-2016, considering the variation of levels pm2.5, under the principal stratification framework. Second, we take advantage of the flexible Bayesian non-parametric mixture model for the post-treatment variable (that is, pm2.5) to understand how these effects can vary within strata. Third, we characterize the different identified groups of individuals.
5.1 Data and Study Design
To answer our research question, we have merged two datasets: (i) one dataset containing the information about the NAAQS designations, as well as air pollution levels and the demographic and socioeconomic characteristics in the counties in the eastern U.S., used in the Zigler et al. (2018)’s analysis (data can be found at Zigler, 2017), and (ii) one dataset containing the information on the age-adjusted mortality rate in these counties available by the Center for Disease Control and Prevention of U.S. (Friede et al., 1993).
The initial dataset from Zigler (2017) is made up of 482 counties in the eastern U.S., where national monitoring networks have detected the concentration of pm2.5. In 2005, the EPA designated as non-attainment of the NAAQS these counties where the average concentration of pm2.5 was above $15\mu g/m^{3}$, or otherwise attainment (note that the revision occurred in 1997, but became effective just in 2005 due to legal disputes). States containing counties designated as non-attainment were required to develop or revise State Implementation Plans (SIP) outlining how a non-attainment area will achieve standards with strategies to reduce ambient concentrations of pm2.5.
As noted in the previous literature, there is a large diversity in local actions in response to the non-attainment designation and in the enactment of subsequent SIPs (Greenstone, 2004; Zigler et al., 2018). Due to this diversity, the direct analysis of the effect of the designation is very similar to an intention-to-treat analysis. This highlights the importance of conducting a principal stratification analysis to compare the effects on health outcomes across different locations: specifically, it contrasts the effects in areas where pollution was effectively reduced by the non-attainment designation (associative negative effects) with those areas where pollution was not measurably affected (dissociative effect), or where it actually increased (associative positive effects).
The study design specifies a baseline period from 2000 to 2005. During this time, each county, $i$, in the dataset is classified according to its attainment status under the EPA 2005 NAAQS revisions. We define a binary treatment variable, $T_{i}$, where $T_{i}=1$ indicates that the i-th county was designated as non-attainment and therefore was required to develop or revise its State Implementation Plans. In contrast, $T_{i}=0$ indicates that the i-th county was designated as attainment and did not have these requirements.
For this period, we also have the average ambient concentration of pm2.5(from the EPA monitoring locations within each county), as well as a number of counfounders such as census variables like: the percentage of Hispanic and black residents; the average household income; the percentage of females; the average house value; the proportion of residents in poverty; the proportion of residents with a high school diploma; the smoking rate; the population; the percentage of residents in urban area; the employment rate (the percentage of the workforce employed); the percentage of the move in the last 5 years. We also have meteorological variables such as the averages of daily temperatures and the relative average of humidity; the dew point (the temperature at which air becomes saturated with moisture, leading to the formation of dew or condensation).
Consistently with Zigler et al. (2018), we identify as the follow-up period a period after 5 years from the implementation of the 2005 NAAQS revision. For this period, we have data on the levels of pm2.5(2010-2012) from Zigler (2017). The levels of pm2.5in the follow-up period are compared with the levels in the baseline period to establish the decrease / increase in air pollution following the implementation of NAAQS. This variable serves as the post-treatment variable $P_{i}$. Furthermore, we gather the publicly available age-adjusted mortality rate of all-cause mortality (2010-2016) from the website of the US Center for Disease Control and Prevention (CDC) (Friede et al., 1993). Mortality rates in the follow-up period are also contrasted with the rates in the baseline period to assess the impact of NAAQS on mortality rates accounting for baseline rates. This variable serves as our outcome variable $Y_{i}$.
The final dataset, obtained by merging the previously mentioned data sources and cleaning the data set to avoid missing data, comprises 384 counties, 270 of which were designed as attainment and 114 as non-attainment. In the Supplementary Material, Figure F.1 maps this final analysis data set.
5.2 Results
We apply our proposed model to the 384 counties in the eastern US described above, including all covariates, census and meteorological variables, in the weights of the post-treatment variable mixture, while for the outcome model we use the linear model in (10). The model identifies the three strata: the dissociative stratum with 115 counties ($30\%$ of the total counties analyzed), the associative positive stratum with 48 counties ($12.5\%$), and the associative negative stratum with 221 counties ($57.5\%$). According to the definition of the three strata and as visualized in the image on the left in Figure 2, the dissociative stratum, identified with the color yellow, is composed of counties where the NAAQS revision does not substantially affect the level of pm2.5 , in fact, the expected value of $\mathbb{E}[P_{i}(1)-P_{i}(0)|S_{i}^{(1)}=S_{i}^{(0)}]$ for the counties allocated to this strata has a median close to zero and the first and third quantiles of $-0.29\mu g/m^{3}$ and $0.07\mu g/m^{3}$. The associative negative stratum, identified with the color green, is made up of counties where the implementation of environmental plans significantly decreases the levels of pm2.5 . Specifically, in these counties, the NAAQS revision reduced by $0.9\mu g/m^{3}$ the median levels pm2.5 , with first and third quantiles of $-1.46\mu g/m^{3}$ and $-0.38\mu g/m^{3}$. The associative positive stratum, identified with the color red, is made up of counties where the revision of NAAQS was associated with increases in pm2.5 levels by $0.8\mu g/m^{3}$ in the median. However, this latter stratum has the highest uncertainty for $\mathbb{E}[P_{i}(1)-P_{i}(0)|S_{i}^{(1)}\prec S_{i}^{(0)}]$, with values of the first and third quantiles equal to $0.13\mu g/m^{3}$ and $1.34\mu g/m^{3}$, respectively.
The corresponding distributions of the expected dissociative/associative effects are reported in the right image in Figure 2 and show the effect of the attainment or non-attainment designations on the mortality rate conditional to the three strata, i.e., conditional to the heterogeneity in the causal effects of the NAAQS revision on the level of pollution. The EDE and the EAE+ have a similar increase in the median $2.7{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$ of the mortality rate when environmental plans are applied. The means of these two effects are $0.8{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$ and $1.3{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$, respectively. For both principal causal effects, the interquartile interval includes the zero value. This means that in counties where the NAAQS revision does not affect the level pm2.5 or where this level increases, the mortality rate is also not affected. The EAE-, in green color on the right image of Figure 2, assumes negative values, indicating that the implementation of environmental plans, in the counties where the regulations affect the level of pm2.5 , significantly reducing it, also reduces the median age adjusted mortality rate. Specifically, the mortality rate decreases by $11.9{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}$ in the median and with an interquartile interval of $[-15.2{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}},-9.1{}^{\text{o}}\mkern-5.0mu/\mkern-3.0mu_{\text{oooo}}]$.
Figure 2: Boxplots of the three identified strata for (left) the conditional average of the difference of the post-treatment variables, and (right) the expected associative/dissociative effects. The light-blue vertical lines show the value zero, identifying the null effects.
In addition, it is our interest to characterize the three different strata (Figure F3 in the Supplementary Materials visualizes the average of observed covariates within each stratum, reported as colored lines, and compares them with the average of covariates between the full 384 counties). The associative positive stratum and the dissociative stratum are composed of rural counties with a smaller population density, where the population has lower income with respect to the mean of the overall counties, and with a small employment rate. In contrast, in countries of the associative positive stratum, the percentage of Hispanics is lower than the average, in favor of a higher percentage of whites and Asians, while the dissociative stratum is characterized by a higher percentage of Hispanics and a smaller community of black people, with a higher level of education. Furthermore, the associative positive stratum is distinguished by a high smoke rate. In contrast, the associative negative stratum identifies urban counties with a high population density. This population is characterized by higher income, but also a higher percentage of poor people and lower levels of education. The meteorological variable, such as the averages of daily temperatures, the relative average of humidity, and the dew point, appears to play an important role in the distinction of strata, that is, in the characterization of the different effects of the implementation of environmental plans on the level of pm2.5. We refer to the Supplementary Materials for more details about the probability of each county to belong in each of the three strata.
6 Discussion
In this paper, we proposed the confounders-aware shared atoms Bayesian mixture model to address critical challenges within the principal stratification framework. Our approach leverages the dependent Dirichlet process to define a highly flexible and adaptable model for the potential post-treatment variables. Our proposed idea of allowing the distribution of potential post-treatment variables to share information through the shared-atoms of the Bayesian nonparametric prior enables the accurate identification of principal strata while relaxing the standard assumptions of the principal stratification causal framework. Our proposed model data-adaptively identifies the principal strata and imputes missing post-treatment variables and outcomes without the need to set predefined (and subjective) thresholds on the continuous post-treatment values to identify the principal strata. Moreover, the proposed model allows for the quantification of the uncertainty in membership in the principal strata. Theorem 1 provides identification for the proposed causal estimands, while Theorem 2 illustrates how the probability of a unit belonging to a dissociative stratum is non-zero and depends on its observed covariates.
The performance of the proposed model is evaluated through an extensive set of Monte Carlo simulations where the performance of our proposed model is compared and contrasted with the performance of the proposed models by Schwartz et al. (2011) and Lu et al. (2023). The results demonstrate that our model consistently outperforms these alternatives, achieving lower bias, better principal strata identification, and accurate estimation of principal causal effects across various scenarios.
The proposed model is used to assess the effectiveness of a previous revision of NAAQS in reducing levels of pm2.5and, in turn, reducing the mortality rate in treated counties. Specifically, we find significant effects on the reduction of mortality rate in areas where pollution was effectively reduced by the non-attainment designation (associative negative effects). In contrast, those areas where pollution was not measurably affected (dissociative effect) or where it actually increased (associative positive effects) did not show any significant effect on the mortality rate.
Future research could build upon our methodology by adopting more flexible approaches to model the final outcome. While our current model uses a linear specification for the outcome, future work could explore non-linear and more sophisticated modeling techniques. In this study, we concentrated on innovations in modeling the post-treatment variable and intentionally kept the outcome model simpler to avoid unnecessary complexity. However, exploring more complex options for outcome models remains an opportunity for further research. Furthermore, our methodology could be adapted to accommodate continuous treatments, similar to the approach suggested by Antonelli et al. (2023), or extended to handle the time-to-event setting as in Xu et al. (2022). We also leave the exploration of these extensions to future studies.
Acknowledgement
Supplementary material
The Supplementary Material includes (A) the explanation of conjugate prior distribution for probit regression parameters and its posterior distribution, (B) the proof Theorem 1 and (C) Theorem 2, (D) the posterior computation, (E) more details of simulations, and (F) more information and results of the application.